### v1

# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

Abstract.
We prove a categorical homological mirror symmetry result for a blow-up of an abelian surface times $\mathbb{C}$, on the complex side. Specifically, the first author’s paper [Can20] provides evidence for homological mirror symmetry for genus 2 curves $\Sigma_{2}$ on the complex side, where $\Sigma_{2}$ is a hypersurface in an abelian surface $V\cong T^{4}$. To obtain a mirror to $\Sigma_{2}$, the generalized SYZ approach [HV00, AAK16] was used. That is, $(\text{Bl}_{\Sigma_{2}\times\{0\}}V\times\mathbb{C}_{y},y)=:(X,y)$ admits a Lagrangian torus fibration which has an SYZ mirror Landau-Ginzburg model $(Y,v_{0})$, known as the generalized SYZ mirror of $\Sigma_{2}$. The mirror to $X$ - without the superpotential - is obtained by removing a generic fiber from $(Y,v_{0})$. In this paper, we prove a categorical homological mirror symmetry result for $X$. To do so, we equip the symplectic mirror with a Fukaya category which involves both partial and full wrapping in the base of a symplectic Landau-Ginzburg model. Semi-orthogonality appears in the categorical invariants on both sides of HMS.
1. Introduction
1.1. Background and motivation
Homological mirror symmetry (HMS) was conjectured by Kontsevich in [Kon95]. It is an equivalence, for certain pairs of complex manifold $X$ and symplectic manifold $Y$, of the bounded derived category of coherent sheaves and the Fukaya category [FOOO09], [Aur14], that is,
(1.1) $D^{b}\text{Coh}(X)\cong D^{\pi}\text{Fuk}(Y).$
The symplectic side is referred to as the “A-side” and the complex side as the “B-side”. So here $X$ is on the B-side and $Y$ is on the A-side. With the structure of a special Lagrangian torus fibration on a symplectic manifold, Strominger-Yau-Zaslow [SYZ96] formulated SYZ mirror symmetry, a geometric way to construct the complex mirror.
To find a mirror to a hypersurface $H$, the generalized SYZ construction can be used [HV00]. This is because a hypersurface doesn’t have an evident Lagrangian torus fibration (the way a toric or abelian variety $V$ does), so we instead consider the intermediary manifold $X$ given by the deformation to its normal cone, because that one does admit such a fibration. In this paper, we prove a categorical HMS result for the mirror $Y$ to the $X$ arising from the hypersurface $H$ given by a 1-parameter family of genus curves in a 4-torus $V$ (principally polarized abelian surface) considered in [Can20].
This overview of the generalized SYZ construction is based on [AAK16]. Let $V$ be a toric variety or a principally polarized abelian variety of complex dimension $n$. Let $\mathcal{L}\to V$ be an ample holomorphic line bundle with section $s$, and $H=s^{-1}(0)$ its divisor, a complex hypersurface of $V$. Let $X:=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}$.
Remark 1.1.
$X$ is known as the deformation to the normal cone because the blow-up map composed with the projection to the $\mathbb{C}$ factor $y:X\to\mathbb{C}$ parametrizes a deformation of a smooth generic fiber to a singular fiber containing the exceptional divisor $E$ over $0\in\mathbb{C}$. Recall that the exceptional divisor is the projectivization of the normal bundle to the submanifold being blown-up, here $\mathbb{P}(\mathcal{N}_{H\times\{0\}\backslash V\times\mathbb{C}})$ where $\mathcal{N}_{H\times\{0\}\backslash V\times\mathbb{C}}\cong(\mathcal{L}\oplus\mathcal{O}_{\mathbb{C}})|_{H\times\{0\}}\to H\times\{0\}$.
The geometry of the blow-up $X$.
(1.2) $X=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}=\{(x_{1},\ldots,x_{n},y,(u_{1}:u_{2}))\in\mathbb{P}(\mathcal{L}\oplus\mathcal{O})\mid s(x_{1},\ldots,x_{n})u_{2}=u_{1}y\}$
$\pi:X\xrightarrow{\text{blow-up }}V\times\mathbb{C},\quad y:X\xrightarrow{\text{pr}_{\mathbb{C}}\circ\pi}\mathbb{C}.$
We can canonically identify the preimage $y^{-1}(t)$ for $0\neq t\in\mathbb{C}$ with $V$ because for ${x}:=(x_{1},\ldots,x_{n})\in V$, then $u_{1}/u_{2}=s({x})/t$ is fixed. As outlined in Remark 1.1, over 0 we have a deformation to $y^{-1}(0)\cong V\cup_{H}E$ intersecting along $H$, as follows, see [AAK16, Proposition 7.3] for more details. The exceptional divisor $E$ is a $\mathbb{P}^{1}$-bundle
$\text{pr}_{V}\circ p:E\to H$
where $p:=\pi|_{E}$, given by
$E=\pi^{-1}(H\times\{0\})\ni({x},s({x})=0,(u_{1}:u_{2}))\mapsto{x}\in H$
for any $(u_{1}:u_{2})\in\mathbb{P}^{1}\cong\mathbb{P}(\mathcal{L}_{{x}}\oplus\mathcal{O}_{\{0\}})$. Also in $y^{-1}(0)$ we have the proper transform $\tilde{V}$ of $V\times\{0\}$, which is
$\tilde{V}=\overline{\pi^{-1}((V^{0}\backslash H)\times\{0\})}=\{({x},0,(1:0))\mid{x}\in V\}\cong V$
where $V^{0}$ is obtained by removing the toric divisors if any (not applicable for an abelian variety). Intersecting these two sets
$\tilde{V}\cap E=\{({x},0,(1:0)\mid{x}\in H\}\cong H.$
Remark 1.2 (Lagrangian torus fibration on $X$ [AAK16, §4]).
The moment map on $V\times\mathbb{C}$ is obtained from the torus action $(e^{i\theta_{1}}x_{1},\ldots,e^{i\theta_{n}}x_{n},e^{i\theta_{n+1}}y)$. Then using cut and paste operations as in [Sym03] on the image of the moment map, the moment polytope, this corresponds to blowing up $V\times\mathbb{C}$ along $H\times\{0\}$. The result is a new polytope that is the base of a Lagrangian torus fibration on $X$.
Generalized SYZ mirror symmetry. Here is a rough outline of the generalized SYZ procedure of [AAK16], which we will later describe in more detail for the setting of this paper. Let
$\Omega_{V\times\mathbb{C}}=i^{n+1}\bigwedge_{j=1}^{n}d\log x_{j}\wedge d\log y$
be a meromorphic $(n+1)$-form on $V\times\mathbb{C}$ with simple poles on the toric divisors $D_{V\times\mathbb{C}}=(V\times 0)\cup(D_{V}\times\mathbb{C})$ where $D_{V}=V\backslash V^{0}$, which is empty if $V$ is an abelian variety. Then we have a canonical form on $X$ given by
$\Omega=p^{*}\Omega_{V\times\mathbb{C}}$
with anticanonical divisor
$D=\tilde{V}\cup p^{-1}(D_{V}\times\mathbb{C}).$
[…]
$\Delta=\{(\xi_{1},\ldots,\xi_{n},\eta)\in\mathbb{R}^{n+1}\mid\eta\geq\text{Trop}(s)(\xi_{1},\ldots,\xi_{n})\}.$
For an abelian variety there are additional steps to quotient by the defining lattice. Both $X$ and $Y$ admit Kähler structures and the holomorphic functions $y:X\to\mathbb{C}$, $W_{0}:Y^{0}\to\mathbb{C}$, and $v_{0}:Y\to\mathbb{C}$ are Landau-Ginzburg models. In this paper, we take $X$ without superpotential to be the complex side (B-side) of mirror symmetry and $(Y^{0},W_{0})$ to be a symplectic Landau-Ginzburg model as defined in [ACLL23b, Definition 1.1].
Example 1.3.
Take $H=\{1\}\subset V=\mathbb{C}^{*}/\mathbb{Z}$ to be a point in an elliptic curve. Then $(Y,-v_{0})$ is the total space of the Tate curve over a disc, of real dimension 4. The structure sheaf on the point $H$ is mirror to the Lagrangian thimble to the critical point of the Tate curve.
Related work. Evidence for categorical HMS of a 1-parameter family of complex genus 2 curves (as a hypersurface in the Jacobian torus) and their generalized SYZ mirror, was proved in [Can20] for a subcategory of the Fukaya category with fiber Lagrangians $\hat{\ell}_{k,[0]}$ and $\omega_{B}=0$ in the notation of [ACLL, §1.2.3 and Equation (2.36)]. SYZ mirror symmetry was proved in [KL19]. HMS for complex hypersurfaces of $(\mathbb{C}^{*})^{n}$ was proved in [AA24] and SYZ mirror symmetry in both directions for hypersurfaces of toric varieties in [AAK16, CLL12]. Another related work includes [Aur18], relating the Fukaya category of LG models of the form $((\mathbb{C}^{*})^{n},f)$ to the wrapped Fukaya category of a smooth fiber and of its complement in $(\mathbb{C}^{*})^{n}$; note that the hypersurface is on the A-side while here it is on the B-side. HMS for toric blow-ups has been studied in [HH22, §6.3], blowing up at toric fixed points, in contrast to hypersurfaces away from toric divisors as here. In the works [VWX23, Ker08] semi-orthogonal decompositions show up in the toric setting, as they do in our abelian variety setting here. Furthermore, [GKR17] discuss other versions of mirror symmetry for general type varieties.
1.2. Statement of main theorem
Generalized SYZ mirror symmetry for $V$ an abelian surface. We describe $X$ and $Y$ in more detail when $V$ is an abelian surface ($n=2$) and $H$ is a genus 2 curve. This is the setting of [Can20]; the HMS result there was for $H$ and $(Y,-v_{0})$. The HMS result here is for $X$ and $(Y^{0},W_{0})$.
Let $V=(\mathbb{C}^{*})^{2}/\tau\mathbb{Z}^{2}$ be the principally polarized abelian surface (topologically a 4-torus $T^{4}$) corresponding to the complex modular parameter $\tau=\frac{i}{2\pi}\log t\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)$, $0<t\in\mathbb{R}_{\ll 1}$, defined by quotienting by
$(\tau n)\cdot(x_{1},x_{2})=(e^{2\pi i\sum_{j=1}^{2}\tau_{1j}n_{j}}x_{1},e^{2\pi i\sum_{j=1}^{2}\tau_{2j}n_{j}}x_{2}),\quad(x_{1},x_{2})\in(\mathbb{C}^{*})^{2}.$
Then $V$ admits a degree 1 ample line bundle $\mathcal{L}=(\mathbb{C}^{*})^{2}\times\mathbb{C}/\tau\mathbb{Z}^{2}$ defined by quotienting by
$(\tau n)\cdot(x_{1},x_{2},v)=(e^{2\pi i\sum_{k=1}^{2}\tau_{1k}n_{k}}x_{1},e^{2\pi i\sum_{k=1}^{2}\tau_{2k}n_{k}}x_{2},e^{-\pi i(n^{T}\tau n)}\prod_{j=1}^{2}x_{j}^{-n_{j}}v)$
with section given by the theta function
$\vartheta(\tau,x):=\sum_{n\in\mathbb{Z}^{2}}e^{\pi in^{T}\tau n}x_{1}^{n_{1}}x_{2}^{n_{2}},\quad x=(x_{1},x_{2})\in(\mathbb{C}^{*})^{2}.$
The 0-set of the theta function in $V$ is a genus 2 curve
$H:=\{x\in V\mid\vartheta(\tau,x)=0\}=\Sigma_{2}\subset V.$
Define $X:=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}$ with blow-up map $\pi:X\to V\times\mathbb{C}$, exceptional divisor $p:E\to H\times\{0\}$, and inclusion $i:H\times\{0\}\to V\times\mathbb{C}$. To obtain the generalized SYZ mirror, we have an additional step of quotienting the above set-up by $\tau\mathbb{Z}^{2}$ after restricting $Y$ to a neighborhood where the group acts freely [Can20, Eqs. (3.10), (3.24), (3.25)]. For small $\varepsilon<1$,
(1.3) $\Delta_{\tilde{Y}}$ $=\{(\xi_{1},\xi_{2},\eta)\in\mathbb{R}^{3}\mid\eta\geq\operatorname{Trop}(\vartheta)(\xi_{1},\xi_{2})\}$
$\Delta_{Y}$ $=(\Delta_{\tilde{Y}})|_{\chi^{(0,0,1)}\leq\varepsilon}/\tau\mathbb{Z}^{2}$
where $\chi^{(0,0,1)}$ is the complex toric coordinate corresponding to the $\eta$ direction, $(0,0,1)$. We will denote it by $v_{0}$ following [Can20, AAK16]. The polytope $\Delta_{\tilde{Y}}$ looks like a bowl tiled by infinitely many hexagons and filled by an amount of $\varepsilon$. Recall from toric geometry that each vertex $v$ corresponds to a chart $\text{Spec}(\mathbb{C}[y_{1}^{v},y_{2}^{v},y_{3}^{v}])\cong\mathbb{C}^{3}$ in local inhomogeneous coordinates $y_{1}^{v},y_{2}^{v},y_{3}^{v}$, and edges define transition functions. The reason this polytope defines the SYZ mirror is because its vertices and edges match the SYZ charts and transition functions. Let
$v_{0}:=y_{1}^{v}y_{2}^{v}y_{3}^{v}.$
Then $(Y,-v_{0})$ is the generalized SYZ mirror to $H$. The relation between $(Y,v_{0})$ and $(Y^{0},W_{0})$ is
(1.4) $W_{0}=T^{\epsilon}(v_{0}-1)$
and
[…]
where $T$ is the Novikov parameter and $\epsilon>0$ is small (different from the $\varepsilon$ in the domain of definition in Equation (1.3)). The Novikov parameter parametrizes the complex structure on $Y$ mirroring symplectic structures on $X$, the way $\tau$ parametrized in the other direction, the complex structure on $X$ mirroring symplectic structures on $Y$. Here $W_{0}:Y^{0}\to\mathbb{C}^{*}$ is a $T^{4}$-fibration over $\mathbb{C}^{*}$ with one singular fiber over $T^{\epsilon}$ (where $v_{0}=0$). We removed a smooth fiber from $Y$, so the singular fiber remains in $Y^{0}$. We use the symplectic form on $Y$ constructed in [Can20, §3.5], restricted to $Y^{0}$. The total space of $Y$ is smooth, and $\omega$ is modeled on the standard form for $\mathbb{C}^{3}$ near critical points of $v_{0}$ (the north and south pole of the banana manifold of three $\mathbb{P}^{1}$’s identified at their poles), while it is modeled on the standard product form for $\mathbb{CP}^{2}(3)\times\mathbb{D}$ away from the critical points.
Remark 1.4.
What was called $x,y,z$ in [Can20, Figure 6, §3.3] is $y_{1}^{(-1,-1,0)},y_{2}^{(-1,-1,0)},y_{3}^{(-1,-1,0)}$ here. And $\tau$ in [Can20] has been renamed $t$ here, in order to reserve $\tau$ for the more standard notation of the complex modular parameter.
To recap, here are the three forms of SYZ mirror symmetry in [AAK16] in the case of $V$ an abelian surface.
Theorem 1.5 ([AAK16, Theorem 10.4]).
  1. (1)
    $Y^{0}$ is SYZ mirror to $X^{0}$.
  2. (2)
    $(Y^{0},W_{0})$ is SYZ mirror to $X$.
  3. (3)
    $(Y,-v_{0})$ is generalized SYZ mirror to $H$.
HMS for the third one is examined in [Can20]. HMS for the second one is the subject of this paper.
Remark 1.6.
In the case of an abelian variety (topologically a compact torus), note that since $W_{0}$ is a symplectic fibration with fibers given by closed, compact tori with volume, we cannot be in the exact setting by Stokes’ theorem. For symplectic form $\omega$, if $\omega|_{W_{0}^{-1}(z)}]=0$ then $\int_{W_{0}^{-1}(z)}\omega^{n}=0$ where $W_{0}^{-1}(z)\cong T^{2n}$, contradiction.
To state our main theorem involving the Fukaya category of $(Y^{0},W_{0})$, we first describe the B-side category.
Semi-orthogonal decomposition for blowups. We adapt the diagram at the start of [Orl92, §4] where their $\tilde{X}\to X$ is our blow up $\pi:\text{Bl}_{H\times\{0\}}V\times\mathbb{C}\to V\times\mathbb{C}$, their $p:\tilde{Y}\to Y$ is the exceptional divisor which is our $\mathbb{P}^{1}$-bundle $p:E\to H\times\{0\}$, and $j:E\to X$ is inclusion. That is, the following commutes:
[…]
Then [Orl92, Assertion 4.2(b), Theorem 4.3] assert that there is a semi-orthogonal decomposition
(1.6) $D^{b}\mathrm{Coh}(X)=\langle j_{*}(p^{*}D^{b}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1)),\pi^{*}D^{b}\mathrm{Coh}(V\times\mathbb{C})\rangle$
where $\mathcal{O}_{p}(-1):=\mathcal{O}_{E}(E)$. In the first factor, this means that a sheaf on $H$ is pulled back to the exceptional divisor and then extended by 0 to all of $X$.
This result holds, roughly speaking, because of Beilinson’s spectral sequence in [Huy06, Proposition 8.28] and the semi-orthogonality of $\mathcal{O}_{\mathbb{P}^{n}}$ with $\mathcal{O}(1)_{\mathbb{P}^{n}}$, namely that $H^{0}(\mathbb{P}^{n},\mathcal{O}(1))\cong\mathbb{C}[x_{0},\ldots,x_{n}]_{1}$ (degree 1 homogeneous polynomials in $(n+1)$-variables) but $H^{0}(\mathbb{P}^{n},\mathcal{O}(-1))=0$.
Main Theorem.
Definition 1.7 (notation $D^{b}_{\mathcal{L}}$, [Can20]).
Let $\text{pr}_{V}:V\times\mathbb{C}\to V$ and $\text{pr}_{\mathbb{C}}:V\times\mathbb{C}\to\mathbb{C}$ be the projections to each factor. We define the following subcategories.
  1. (1)
    $D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$: full subcategory of objects $\{\mathcal{L}^{\otimes j}[k]\}_{j,k\in\mathbb{Z}}$
  2. (2)
    $D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$: full subcategory of objects $\{(\text{pr}_{V}^{*}\mathcal{L}^{\otimes j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}})[k]\}_{j,k\in\mathbb{Z}}$
  3. (3)
    $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$: full subcategory of objects $\{\mathcal{L}^{\otimes j}[k]|_{H}\}_{j,k\in\mathbb{Z}}$
  4. (4)
    $D^{b}_{\mathcal{L}}\mathrm{Coh}(X)$: full subcategory of objects in $j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{E}(E))$ and $\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, and including shifts $[k]$
Theorem 1.8.
There exist two full subcategories $\mathcal{F}_{1}$ (U-shaped Lagrangians) and $\mathcal{F}_{2}$ (Lagrangians fibered over cotangent fibers) of the Fukaya category $H^{*}FS(Y^{0},W_{0})$ which are equivalent to $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ and $D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, respectively. Furthermore, for the full subcategory $\mathcal{F}:=\mathcal{F}_{1}\cup\mathcal{F}_{2}$, there is an equivalence
(1.7) $\mathcal{F}\xrightarrow[]{\cong}D^{b}_{\mathcal{L}}\mathrm{Coh}(X).$
In particular, the semi-orthogonality in $D^{b}_{\mathcal{L}}Coh(X)$ is respected.
Remark 1.9.
The wrapping phenomenon that appears when we puncture $Y$ to $Y^{0}$ also appears in related settings such as [GPS24, §3.3] where stop removal corresponds to localization.
Remark 1.10.
For a suitable definition of $FS(Y^{0},W_{0})$, upgrading the theorem to an $A_{\infty}$-result, and for a suitable adaptation of core HMS [PS23], we expect this theorem can be upgraded to the originally conjectured equivalence in Equation (1.1) which would be $D^{b}Coh(X)\cong D^{\pi}FS(Y^{0},W_{0})$.
Remark 1.11.
We directly construct Lagrangians and calculate their morphisms in $(Y^{0},W_{0})$. Another approach could be to combine the left diagram of Figure 1 for $H^{0}FS(Y,v_{0})$ from [Can20], prove a similar diagram for $H^{0}FS(V^{\vee}\times\mathbb{C}^{*}_{y^{\vee}},y^{\vee})$ as in the right diagram of Figure 1, and then combine them as suggested in [Aur18, Remark 1.5] and also discussed in [Syl19, §3.3: Gluing].
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}\mathrm{FS}(Y,v_{0})$$\scriptstyle{\iota^{*}}$$\scriptstyle{\text{HMS on }V}$$\scriptstyle{\text{HMS on }H=\Sigma_{2}}\scriptstyle{\cup}$
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}\mathrm{FS}(V^{\vee}\times\mathbb{C}^{*}_{y^{\vee}},y^{\vee})$$\scriptstyle{\mathrm{pr}_{V}^{*}}$$\scriptstyle{\text{HMS on }V}$
Figure 1. HMS squares for two components of Remark 1.11
Outline of the paper. In Section 2 we describe the B-side semi-orthogonality of $D^{b}\text{Coh}(X)$. In Section 3, we define the mirror Lagrangian objects in Subsection 3.1 and morphisms in the Fukaya category of $(Y^{0},W_{0})$ in Subsection 3.2. This uses standard theory where the details to fill in for this setting are to define the Lagrangians, continuation maps, and prove the latter are well-defined. We define how Lagrangians wrap in Subsection 3.2.1, and give an overview of how to pass from the cochain complex to the localized category in 3.2.2. Then in Subsections 3.2.3 and 3.2.4 we define the quasi-units we localize at, and their inverses, the continuation maps. Finally in Subsection 3.2.5 we define the morphisms by taking a homotopy colimit over multiplication by quasi-units, using a model from [AA24] that amounts to wrapping one way on the input Lagrangian and in the opposite direction on the output in Equation (3.29). In Subsection 3.2.6 we carry out the calculations of the morphisms in Lemmas 3.20 (within $\mathcal{F}_{1}$), 3.23 (nonzero from $\mathcal{F}_{1}$ to $\mathcal{F}_{2}$), 3.24 (zero from $\mathcal{F}_{2}$ to $\mathcal{F}_{1}$), and 3.25 (within $\mathcal{F}_{2}$). We then move on to the product structure, to prove the HMS map is indeed a functor, in Section 4. This requires an analysis of gradings to determine what intersection points appear in the product, which is Lemma 4.3. We then prove the main theorem, Theorem 1.8, in Section 5.
Acknowledgements. We thank Denis Auroux for helpful comments in formulating the direction of this project and for explaining his joint work [AA24]. We thank Hiro Lee Tanaka, Sheel Ganatra, and Zack Sylvan for helpful discussions on wrapping in Fukaya categories, Christian Schnell and Alexander Kuznetsov for helpful discussions on semi-orthogonal decompositions of $D^{b}\text{Coh}$, and Katrin Wehrheim for helpful discussions on moduli spaces. Thanks to Haniya Azam, Hiro Lee Tanaka, and Sheel Ganatra for helpful comments on an earlier draft. C. Cannizzo was partially supported by NSF DMS-2316538. She also thanks the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures (GAUS) at Goethe Universität-Frankfurt for hosting her during the completion of this work.
2. B-side semi-orthogonal decomposition
Let $\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}}:=\text{pr}_{V}^{*}\mathcal{L}^{j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}}$. Morphisms in the bounded derived category of coherent sheaves are Ext groups. Recall Equation (1.6) that
$D^{b}\mathrm{Coh}(X)=\langle j_{*}(p^{*}D^{b}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1)),\pi^{*}D^{b}\mathrm{Coh}(V\times\mathbb{C})\rangle$
where
[…]
commutes.
2.1. Orthogonality: morphisms from $H$ to $V\times\mathbb{C}$
By semi-orthogonality, we know that for $k_{1},k_{2}\in\mathbb{Z}$,
(2.1) $\operatorname{Ext}(\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}}),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{p}(-1)))=0$
which we will see matches with the A-side. We also need to show that morphisms in the other direction match with the A-side, so we consider those next.
2.2. Morphisms in the other direction: $V\times\mathbb{C}$ to $H$
By the adjunction formula $K_{X}=\pi^{*}K_{V\times\mathbb{C}}\otimes\mathcal{O}(E)$. As in [Huy05, Proposition 2.5.5], we define the relative canonical bundle as follows,
$K_{E/X}:=K_{E}\otimes j^{*}K^{-1}_{X}=j^{*}(K_{X}\otimes\mathcal{O}(E))\otimes j^{*}K^{-1}_{X}=\mathcal{O}_{E}(E)$
where $\mathcal{O}_{E}(E):=j^{*}\mathcal{O}(E)=\mathcal{O}_{p}(-1)$. Furthermore, $j^{!}$ admits a left adjoint $j_{*}$ for $j:E\to X$ where
$j^{!}(\cdot)=j^{*}(\cdot)\otimes K_{E/X}[\dim_{\mathbb{C}}E-\dim_{\mathbb{C}}X]=j^{*}(\cdot)\otimes\mathcal{O}_{E}(E)[-1].$
Thus
(2.2) $\operatorname{Ext}_{X}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
$=\operatorname{Ext}_{E}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E),j^{!}\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
$=\operatorname{Ext}_{E}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E),j^{*}\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)[-1]).$
Note that $\otimes\mathcal{O}_{E}(E)$ is an auto-equivalence of $D^{b}\text{Coh}(X)$, thus its contribution cancels. Also, the statement that Equation (1.5) commutes is that $\pi\circ j=i\circ p$. Thirdly, $p:E\to H\times\{0\}$ is surjective so $p^{*}$ is fully faithful, namely, it is an isomorphism on hom sets. Hence we can simplify to obtain
(2.3) $\operatorname{Ext}_{E}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}),j^{*}\pi^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})[-1])$
$=\operatorname{Ext}_{E}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}),p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})[-1])$
$\cong\operatorname{Ext}_{H}(i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}),i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})[-1])$
$\cong\operatorname{Ext}_{H}(\mathcal{L}^{k_{2}}|_{H},\mathcal{L}^{k_{1}}|_{H})[-1]$
where in the last step we restrict to $H\times\{0\}$. Now we show this matches with the symplectic side, including the shift [1][-1] which corresponds to a choice of Lagrangian grading on the mirror.
3. A-side semi-orthogonality
We define the objects and morphisms in the wrapped derived Fukaya category for the symplectic Landau-Ginzburg model $(Y^{0},W_{0})$ with symplectic form $\omega$ given by that of [Can20, §3.5] restricted to $Y^{0}$. This builds on the constructions of many Fukaya categories: the original $A_{\infty}$-category [FOOO09] and Fukaya-Seidel category for Lefschetz fibrations [Sei08], as well as [AS10, GPS20] for exact symplectic manifolds, and [AA24] for Landau-Ginzburg models, to name a few.
The Fukaya category roughly consists of Lagrangian submanifolds as objects and cochain complexes generated by their intersection points as morphisms, graded by the Maslov index. The differential counts bigons between input and output intersection points, weighted by area. An account of the Fukaya category for arbitrary symplectic form, including non-exact and with B-field, as well as how structure maps change under Lagrangian isotopy, is given in [ACLL23a]. Here we take the B-field to be zero. For gradings in a general symplectic Landau-Ginzburg model including non-exact and non-Lefschetz, we use [ACLL23b, §3].
Since $Y^{0}$ is non-compact, we take the Fukaya-Seidel category on $W_{0}:Y^{0}\to\mathbb{C}$. To use the machinery for Fukaya-Seidel categories, we need to define the continuation maps and show they are well-defined by analyzing the geometry of $(Y^{0},W_{0})$. On the punctured Landau-Ginzburg model, the projection of Lagrangians in $\mathcal{F}_{2}$ under $W_{0}$ is a curve that can fully wrap in the base around the puncture. All projected Lagrangians still partially wrap at infinity, as is defined for Lagrangians in a Landau-Ginzburg model to $\mathbb{C}$ [AS19, Appendix A], [GPS20, §1.2, Example 1], [AA24]. Since fibers and their Lagrangians are compact, we do not wrap Lagrangians in the fibers. These features give the definition of our continuation maps in Definition 3.9.
Remark 3.1 (Two pictorial representations for the base: $\mathbb{C}^{*}$ and cylinder).
The base of $W_{0}$ is topologically a cylinder $\mathbb{C}^{*}$, and a generic fiber is $T^{4}$. Curves drawn in a plane represent projections of Lagrangians under $W_{0}$, while curves drawn on a cylinder represent projections of Lagrangians under $W_{0}$ composed with $\mathbb{C}^{*}\to\mathbb{R}\times S^{1}$ given by $re^{2\pi i\theta}\mapsto(r,\theta)$.
0
$\infty$
x
$L_{i}$
$K_{j}$
Figure 2. Projections of the two types of Lagrangians to the cylinder model for the base of $W_{0}:Y^{0}\to\mathbb{C}$
3.1. Objects in $\mathcal{F}_{1}$, $\mathcal{F}_{2}$
Lagrangians are “thimble-like” in the sense of [GPS24, Definition 8.19], as follows.
Definition 3.2 (Parallel transport, [MS98, Equation (6.2)]).
Let $\Phi_{\gamma(0)\to\gamma(t)}:W_{0}^{-1}(\gamma(0))\to W_{0}^{-1}(\gamma(t))$ denote parallel transport in $(Y^{0},W_{0})$ with respect to $\omega|_{Y^{0}}$ (for $\omega$ on $Y$ defined in [Can20, §3.5]) over an embedded path $\gamma:[0,t]\to\mathbb{C}^{*}$ avoiding the critical value $\text{critv}(W_{0})=-T^{\epsilon}$.
We define two subcategories $\mathcal{F}_{1},\mathcal{F}_{2}$ of the Fukaya-Seidel category of $(Y^{0},W_{0})$. Define a point
(3.1) $b_{S}=-S\in\mathbb{C}^{*},\quad S\in(T^{\epsilon},T^{\varepsilon}).$
[…]
Definition 3.3 (Fiber Lagrangians).
Define slope $k$ Lagrangians in the fiber $W_{0}^{-1}(b_{S})\cong T^{4}$ by
$\ell_{k}:=\{(\xi_{1},\xi_{2},\theta_{1},\theta_{2})\in T^{4}\mid(\theta_{1},\theta_{2})=-k\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)^{-1}\left(\begin{matrix}\xi_{1}\\ \xi_{2}\end{matrix}\right)\}$
where $(\theta_{1},\theta_{2})=\frac{1}{2\pi}(\arg(t_{1}),\arg(t_{2}))$ for $(t_{1},t_{2})\in(\mathbb{C}^{*})^{2}$ the affine coordinates corresponding to monomials $\chi^{(1,0,0)}$ and $\chi^{(0,1,0)}$ on the polytope $\Delta_{Y}$ of Equation (1.3).
Definition 3.4 (Objects of $\mathcal{F}_{1}$).
Objects of $\mathcal{F}_{1}$ are $\{L_{j}[k]\}_{j,k\in\mathbb{Z}}$ defined as
$L_{j}:=\bigcup_{-\infty<t<\infty}\Phi_{\gamma_{L_{j}}(0)\to\gamma_{L_{j}}(t)}(\ell_{j})$
where
$\gamma_{L_{j}}(t)=r(t)e^{2\pi\mathbf{i}\theta(t)}:(-\infty,\infty)\to\mathbb{C}^{*}$
such that
$r(0)=S,\qquad\theta(0)=-\frac{1}{2},\qquad r^{\prime}(t)=1\;\forall t>>0\text{ and }\forall t<<0,\qquad\theta(t)=\begin{cases}\theta_{h_{1}}(L_{j})&\forall t>0\,s.t.~r(t)>R_{4}\\ \theta_{h_{2}}(L_{j})&\forall t<0\,s.t.~r(t)>R_{4}\end{cases},$
for $R_{4}$ as in Equation (3.3) and where $r^{\prime}(t)=1$ by the time $r(t)$ reaches $R_{4}$, so that $\gamma_{L_{j}}$ is a U-shape in the base of $W_{0}$ around $T^{\epsilon}$ whose ends are angular lines at constant angles $\theta_{h_{1}}(L_{j})<\theta_{h_{2}}(L_{j})$ in the interval (1/2,1/2)(-1/2,1/2) on each end. In particular, $\gamma_{L_{j}}(0)=-S$.
Remark 3.5.
U-shaped Lagrangians have appeared in the literature such as [AS19, §A.2], [ACLL23b], and [AA24]. Here we use the model that ends of the U-shape project to angular rays away from 0.
Definition 3.6 (Objects of $\mathcal{F}_{2}$).
Objects of $\mathcal{F}_{2}$ are $\{K_{j}[k]\}_{j,k\in\mathbb{Z}}$ defined as
$K_{j}:=\bigcup_{1<t<\infty}\Phi_{\delta_{K_{j}}(0)\to\delta_{K_{j}}(t)}(\ell_{j})$
where
$\delta_{K_{j}}(t)=r(t)e^{2\pi\mathbf{i}\theta(t)}:[0,\infty)\to\mathbb{C}$
such that
$r(0)=S,\qquad r(1)=0,\qquad\theta(0)=-\frac{1}{2},\qquad(r(t),\theta(t))=(t,\theta_{h}(K_{j})),\;\forall t>1$
so that $\delta_{K_{j}}$ is an arc from S-S, passing through 0, and is then a single arc at fixed angle $\theta_{h}(K_{j})$. To be the correct mirror, we take the angle to lie above any U-shape ends so $\theta_{h}(K_{j})>\theta_{h^{\prime}}(L_{j^{\prime}})$ in (1/2,1/2)(-1/2,1/2) $\forall j^{\prime}\in\mathbb{Z}$.
Parallel transport starts from the fiber over $\delta_{K_{j}}(0)=-S$, but the Lagrangian is only over the part of the arc for $t>1$ because in $Y^{0}$ we puncture the base at 0 which is where the curve passes through at $t=1$. In other words, $W_{0}(K_{j})$ in the cylinder model is a cotangent fiber of $T^{*}S^{1}$ at height $\theta=\theta_{h}(K_{j})$.
We will define the gradings on the Lagrangians in Subsection 4.1.
3.2. Morphism groups
3.2.1. Wrapping conventions in the cochain complex
For two transversely intersecting Lagrangians $L_{0},L_{1}$, the Floer cochain complex is
(3.2) $CF^{m}(L_{0},L_{1})=\bigoplus_{\begin{subarray}{c}p\in L_{0}\pitchfork L_{1}\\ \deg(p)=m\end{subarray}}\mathbb{C}\cdot p.$
If they are not transverse we perturb by isotopies. Let $L$ be a Lagrangian of the form $L_{j}$ or $K_{j}$. Recall that a Lagrangian isotopy is a homotopy
$\phi:L\times[0,1]\to Y^{0}$
which is a smooth embedding. We have a horizontal distribution
$\mathrm{Hor}_{p}=\left(\ker d_{p}W_{0}\right)^{\perp_{\omega}}=(T_{p}Y^{0}_{W_{0}(p)})^{\perp_{\omega}}\subset T_{p}Y^{0}$
for the symplectic form $\omega$ because $W_{0}$ is a symplectic fibration, that is, $\omega$ restricts to a symplectic form on smooth fibers. Isotopies in a Landau-Ginzburg model [ACLL23b, §2] are constructed by parallel transport over homotopies $\delta(s,t)$ of curves in the base $\mathbb{C}$. Here they partially rotate the angular ends going to infinity, not crossing $\pi$, and fully rotate around 0 in the other direction.
Let $\rho^{t}$ be the flow of the horizontal lift of the Hamiltonian vector field for the Hamiltonian $H$ defined as follows. Let $w$ be the coordinate on the base $\mathbb{C}^{*}$ of $W_{0}$. For $|w|\neq 0$ and $\chi(\arg w):[-\pi,\pi)\to[0,1]$ a bump function supported away from $\pi$, pick $0<R_{1}<R_{2}<T^{\epsilon}<S<R_{3}<R_{4}<T^{\varepsilon}$ and define
(3.3) $H(w):=\begin{cases}\chi(\arg w)\cdot|w|&R_{4}<|w|<T^{\varepsilon}\\ \text{interpolate}&|w|\in[R_{3},R_{4}]\\ 0&R_{2}<|w|<R_{3}\\ \text{interpolate}&|w|\in[R_{1},R_{2}]\\ |w|^{2}&|w|<R_{1}\end{cases}$
where recall $|W_{0}|:Y_{0}\to[0,T^{\varepsilon}]$ arises as $W_{0}$ is a fibration only over a small disc. So within this domain but outside of the annulus
(3.4) $A=\{w\in\mathbb{C}\mid R_{1}<|w|<R_{4}\}$
Lagrangians project under $W_{0}$ to arcs which wrap slightly on one end and fully around 0. Within the smaller annulus $A=\{w\in\mathbb{C}\mid R_{2}<|w|<R_{3}\}$ there is no movement, including at S-S where Lagrangians are defined as parallel transport from in Equation (3.1). Define
$\rho^{t}:Y^{0}\to Y^{0}$
to be the flow of a vector field $\xi^{\#}$ which is the horizontal lift of $\xi$ defined by $\iota_{\xi}\omega_{\mathbb{C}}=dH$ for $\omega_{\mathbb{C}}=\frac{i}{2}dw\wedge d\overline{w}=dw_{x}\wedge dw_{y}=rdr\wedge d\theta=e^{2u}du\wedge d\theta$ for $w=w_{x}+iw_{y}=re^{i\theta}=e^{u+i\theta}$. (Note that $\xi$ is sometimes called $X_{H}$ in literature where $\iota_{X_{H}}\omega_{\mathbb{C}}=-dH$, but in [AA24] $X_{H}$ is used for a wrapping Hamiltonian $H$ in the fiber while $\xi$ realizes the wrapping in the base. Since we are wrapping in the base and not the fiber, to be consistent with them, we use $\xi$.) By [ACLL23b, Remark 3.2], $\omega_{\mathbb{C}}$ is related to $\omega$ by
(3.5) $({W_{0}}^{*}\omega_{\mathbb{C}})|_{\text{Hor}}=|dW_{0}|^{2}_{J,\omega}\omega|_{\text{Hor}}.$
The relation between positively or negatively wrapping with respect to $\lambda:=\frac{1}{2}r^{2}d\theta$ (where $\omega_{\mathbb{C}}=d\lambda$) and counterclockwise/clockwise wrapping around 0 and infinity is as follows.
  1. (1)
    Near $\infty$, by which we mean in the region $R_{4}<r<T^{\varepsilon}$ near the boundary of $W_{0}(Y^{0})$,
    (3.6) $\iota_{-\chi^{\prime}\partial_{r}+\frac{\chi}{r}\partial_{\theta}}(rdr\wedge d\theta)=-dH=-d(\chi(\theta)r)$
    so on $\xi|_{\{R_{4}<r<T^{\varepsilon}\}}=-\chi^{\prime}\partial_{r}+\frac{\chi}{r}\partial_{\theta}$ we find that
    (3.7) $\lambda(-\chi^{\prime}\partial_{r}+\frac{\chi}{r}\partial_{\theta})=\frac{1}{2}r^{2}d\theta(\frac{\chi}{r}\partial_{\theta})=\frac{1}{2}\chi r>0$
    and wrapping is positive and counterclockwise.
  2. (2)
    Around 0 we flow clockwise by $\xi=2\frac{\partial}{\partial\theta}$ as we approach 0. This is because under the conformal coordinate change $w\mapsto\tilde{w}:=1/w$ on $\mathbb{C}^{*}$, then
    (3.8) $\omega_{\mathbb{C}}=\frac{1}{\tilde{r}^{3}}d\tilde{r}\wedge d\tilde{\theta}$
    and for $\tilde{r}>1/R_{1}$
    (3.9) $dH=\frac{-2}{\tilde{r}^{3}}d\tilde{r}=\iota_{-2\partial/\partial\tilde{\theta}}\frac{1}{\tilde{r}^{3}}d\tilde{r}\wedge d\tilde{\theta}$
    while
    (3.10) $\lambda(-2\frac{\partial}{\partial\tilde{\theta}})=-\frac{1}{2\tilde{r}^{2}}d\tilde{\theta}(-2\frac{\partial}{\partial\tilde{\theta}})=\frac{1}{\tilde{r}^{2}}>0$
    so we wrap counterlockwise with increasing $\tilde{r}$. Inverting to $w$, we wrap clockwise with decreasing $r$.
Now let $L^{k}:=\rho^{-\tilde{\epsilon}k}(L)$ for small positive number $\tilde{\epsilon}$. So larger $k$ means more negatively wrapped. Following [AA24, Equation (3.40)], we formally define a directed category $\mathcal{O}$ with objects the Lagrangians $\{L^{j}\}_{j\in\mathbb{Z}}$ so they form an exceptional collection
$\mathcal{O}(L^{j_{1}},L^{j_{2}}):=CF(L^{j_{1}},L^{j_{2}}),\quad j_{1}<j_{2}.$
In other words, when the ends of the input projected Lagrangian near infinity lie above/counterclockwise from the ends of the output, there is a Floer complex over their intersections. In the other direction it is 0.
The differential $\partial:CF^{m}(L_{0},L_{1})\to CF^{m+1}(L_{0},L_{1})$ on the cochain complex of transversely intersecting Lagrangians is
(3.11) $\partial(p)=\sum_{q\in L_{0}\cap L_{1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in H_{2}(Y^{0},L_{0}\cup L_{1})\\ \text{ind}(u)=1\end{subarray}}\#\mathcal{M}(p,q;[u],J)/(u\sim u_{a})e^{-2\pi\int_{\beta}\omega}q$
where
(3.12) $\mathcal{M}(p,q;[u],J):=\left\{u:(\mathbb{R}\times[0,1],\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\})\to(Y^{0},L_{0}\cup L_{1})\middle|\begin{array}[]{l}\overline{\partial}_{J}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{0},\ u(\mathbb{R},1)\subset L_{1},\\ \lim_{s\to-\infty}u(s,t)=q,\\ \ \lim_{s\to\infty}u(s,t)=p\end{array}\right\}$
where we quotient by $u_{a}(s,t)=u(s-a,t)$ for each $a\in\mathbb{R}$, the translation in the $s$ direction, and where $J$ is an almost complex structure compatible with $\omega$ which is also regular, meaning the linearization of $\overline{\partial}_{J}$ is surjective for all $u\in\mathcal{M}(p,q;[u],J)$. In order for this to be well-defined, meaning in order for us to be able to count the moduli space, it suffices for it to be a 0-dimensional compact smooth manifold. The 0-dimensional part is ensured by choosing $p$ and $q$ so that $\deg(p)+1=\deg(q)$. The compact part is ensured by the curves $u$ satisfying a maximum principle, the way holomorphic curves do, and by excluding bubbling. And the smooth manifold part can be ensured by the existence of regular $J$.
This was proven in [Can20] in the setting for $(Y,-v_{0})$. The difference between the set-up in [Can20] and here is that we have a puncture at 0 in the base so therefore we have additional Lagrangians fibered over single arcs that can wind infinitely many times around 0 in the base. So we have several cases to consider.
3.2.2. Algorithm to derive the category
The morphisms are given by the wrapped Lagrangian Floer cohomology theory. They are defined by the property that Hamiltonian-isotopic Lagrangians should be quasi-isomorphic; they should be isomorphic when we derive the category. There is a convenient tool from homotopy theory that does the job, called a homotopy colimit [AS10, AA24]. As explained in the introduction to [Hir03], a goal of homotopy theory is to identify homotopic maps, but a homotopy on topological spaces may not make sense in other categories. To identify homotopic morphisms $f$ and $g$, we define a set called the weak equivalences to be where fgf-g lands in, and then localize the category at weak equivalences. To take a homotopy colimit of a diagram one needs a model category: a category with three sets of morphisms called weak equivalences, fibrations, and cofibrations. Quillen axiomatized such a category as a place where one can do homotopy theory, calling them “a category of models for a homotopy theory” [Qui67]. Note that more modern approaches in homotopy theory use $\infty$-categories instead of model categories. In the wrapped Fukaya category, weak equivalences will be multiplication by quasi-units. Here is the algorithm used to invert the quasi-units:
  1. (1)
    For chain complexes bounded below, over an abelian category with enough projectives, there is a model category structure where the weak equivalences are the quasi-isomorphisms, the fibrations are the epimorphisms, and the cofibrations are injective maps $i:C_{\bullet}\to D_{\bullet}$} where the cokernel complex ${\text{Coker}}(i)_{\bullet}$} is a complex of projective objects [Qui67, §1.2 Example B]. It follows that the cofibrant objects are the complexes whose objects are all projective. In our case ([AA24, Equation (3.40)], [Can20, Equation (4.2)]) the model category will be a category of modules over another category $\mathcal{O}$,
    (3.13) $\mathcal{O}^{*}(L_{0}^{k_{0}},L_{1}^{k_{1}})=\begin{cases}CF^{*}(L_{0}^{k_{0}},L_{1}^{k_{1}})&k_{0}<k_{1}\\ \mathbb{C}\cdot\text{id}&k_{0}=k_{1},L_{0}=L_{1}\\ 0&\text{otherwise}\end{cases}$
    where id is the identity morphism when composed with any $CF^{*}(L_{0}^{k_{0}},L_{1}^{k_{1}})$. Then the model category $\mathcal{A}$ is the category of $\mathcal{O}$-modules with objects $\mathcal{Y}_{L}$ given by Yoneda modules $X\mapsto\mathcal{O}(X,L)$ and quasi-isomorphisms are the quasi-units. Counts of curves between multiple intersection points define the $A_{\infty}$-structure $\mu^{k}$.
  2. (2)
    Construct the homotopy category by deriving $\mathcal{A}$ so that quasi-isomorphisms become isomorphisms. An object is therefore an equivalence class of a chain complex, up to quasi-isomorphism, and a morphism is an equivalence of a chain map, up to chain homotopy. This is a localization process [Wei94, §10.3] which has an $A_{\infty}$-version in [LO06].
    1. (a)
      The directed category $\mathbb{Z}$ indexes the diagram for each Lagrangian $L$, given by $\ldots\to\mathcal{Y}_{L^{k}}\xrightarrow[]{\mu^{2}(e_{L^{k}},-)}\mathcal{Y}_{L^{k+1}}\to\ldots$ for $k\in\mathbb{Z}$ given by multiplication by quasi-units. Define
      $\mathcal{Y}_{L^{\infty}}:=\text{hocolim}(\mathcal{Y}_{L^{k}}).$
    2. (b)
      To compute homs with $\mathcal{Y}_{L^{\infty}}$ we use the standard model in wrapped Floer theory
      $\mathcal{Y}_{L^{\infty}}=\text{Cone}\left(\bigoplus_{k=0}^{\infty}\mathcal{Y}_{L^{k}}\xrightarrow[]{\text{id}-e_{L^{k}}}\bigoplus_{k=0}^{\infty}\mathcal{Y}_{L^{k}}\right).$
Remark 3.7.
The homotopy direct limit over a collection of isotopies avoids the definition depending on a given isotopy. Note that using a quadratic Hamiltonian to wrap as in [Aur14, §4.2] is equivalent to a colimit by linear Hamiltonians (the original definition of the wrapped Fukaya category [AS10, §5.1]) by [OT24, Proposition 6.1.9].
3.2.3. Quasi-isomorphisms: quasi-units
Following the convention of [AA24, Equation (3.42)], a quasi-unit counts certain curves $\Sigma:=\mathbb{D}\backslash\{1\}\to Y^{0}$ with one moving Lagrangian boundary condition $\rho^{-\tilde{\epsilon}g_{e}(s)}(L)$ for $g_{e}:\mathbb{R}\cong\partial\Sigma\to[k,k+1]$ monotonically increasing smooth and constant near the ends, meaning there is some $K\in\mathbb{R}$ so
(3.14) $g_{e}(s)=\begin{cases}k,&s<-K\\ k+1,&s>K\\ \text{interpolate increasing},&s\in[-K,K]\end{cases}$
so the Lagrangian isotopy is traversed backwards along the boundary.
For the PSS isomorphism $PSS:H^{*}(L)\to HF^{*}(L^{k},L^{k+1})$, the quasi-unit is the image of 1, and can be constructed on the chain-level by a disc count,
(3.15) $e_{L^{k}}:=PSS(1)=\left[\sum_{p\in L^{k}\cap L^{k+1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in H_{2}(Y^{0},\bigcup_{s\in\mathbb{R}}\rho^{-\tilde{\epsilon}g_{e}(s)}(L^{k}))\\ \text{ind}(u)=0\end{subarray}}\#\mathcal{M}(p;[u],J)e^{-2\pi\int_{\beta}\omega}\cdot p\in CF(L^{k},L^{k+1})\right]\\$
[…]
$u(s,0)\in\rho^{-\tilde{\epsilon}g_{e}(s)}(L^{k}),\lim_{s\to\pm\infty}u(s,0)=p\}$
[…]
3.2.4. Inverses: continuation maps
Since we will see in Claim 3.18 that the input and output in morphisms wrap in opposite directions, we need an inverse to quasi-units to connect Floer complexes in Lemma 3.14. These will be the continuation maps. The algorithm to define and use them is as follows.
  1. (1)
    Define the isotopies in the moving Lagrangian boundary condition. We use those induced by Equation (3.3).
    1. (a)
      Absorb the moving Lagrangian boundary condition into a perturbation to the Cauchy-Riemann equation [Aur14, Remark 1.10], known as Floer’s equation. Here it is
      (3.17) $(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$
      for a suitable $\tau$ depending on two parameters. Continuation maps count these solutions.
    2. (b)
      Rescale the curve to obtain a solution to the Cauchy-Riemann equation.
  2. (2)
    Deduce results for elliptic operators which are needed to show that the continuation map moduli space in Equation (3.23) can be counted.
    1. (a)
      A 0-dimensional manifold: A regular $J$ exists, meaning the linearization of the Cauchy-Riemann operator at each $u$ in the moduli space is surjective. This is a standard argument provided curves are shown to have a dense set of somewhere injective points.
    2. (b)
      Compact: this follows from Gromov compactness. This is a standard argument provided there’s a maximum principle that the curves in the moduli space stay in a compact subset for a given homology class $[u]$.
We follow the methodology of [AA24, Chapter 3], noting that we have no $X_{H}$ term since we have no wrapping in the fiber. Another related setting is that of [GPS24, §8.6], since curves projected to the base are then in an exact symplectic manifold.
[…]
Definition 3.8.
The moving Lagrangian boundary used to define the continuation maps is
(3.18) $\Lambda_{s}(L)=\rho^{-\tilde{\epsilon}\tilde{\chi}(s)}(L),\quad s\in\mathbb{R}$
where $\tilde{\chi}:\mathbb{R}\to[0,1]$ is a smooth monotonically decreasing function, constant near -$\infty$ and $\infty$ at 1 and 0 respectively, and $\tilde{\epsilon}$ is small.
This isotopy moves in the opposite direction of the moving boundary condition on the quasi-unit. Because $L^{j}=\rho^{-\tilde{\epsilon}j}(L)$, the isotopy $\Lambda_{s}$ moves $L^{j+1}$ to $L^{j}$. That is, there exists $\tilde{K}$ so that
(3.19) $\Lambda_{s}(L^{j})=\begin{cases}L^{j+1},&s<-\tilde{K}\\ L^{j},&s>\tilde{K}\end{cases}$.
So in words, $\Lambda_{s}(L):\mathbb{R}\to Y^{0}$ rotates Lagrangian ends of $W_{0}(L)$ counterclockwise outside a compact set around 0, meaning $\frac{d}{ds}\theta_{h}(\Lambda_{s}(L))>0$ for $\theta_{h}$ as in Definitions 3.4 and 3.6. When $L=K_{j}$, $W_{0}(\Lambda_{s}(L))$ also fully wraps clockwise around 0. We cancel this movement by introducing a perturbation term moving in the direction of $\rho^{\tilde{\epsilon}\tilde{\chi}(s)}$ on the boundary:
(3.20) $(\xi^{\tau})^{\#}\otimes d\tau$
for $\tau(s,t):\mathbb{R}\times[0,1]\to\mathbb{R}$ a smooth function such that
(3.21) $\tau(s,j)=\tilde{\epsilon}\tilde{\chi}(s),\quad j=0,1.$
Thus $\tau$ is a way to extend the Lagrangian perturbation on the boundary of the disc to the whole disc, providing enough parameters in the perturbation to ensure transversality, while also canceling the moving Lagrangian boundary condition in $\Lambda_{s}(L)$. It is valued in the vector $(\xi^{\tau})^{\#}$ on the target and $d\tau$ inputs vectors on the domain, so tensored together it is the same type as $du$. It describes how $du$ should be perturbed infinitesimally over the domain of $u$.
To ensure enough perturbation parameters to prevent solutions to the Floer equation that have a lot of symmetry because they are nowhere somewhere injective, we need the full set of two parameters on the domain $\Sigma$ in both $s$ and $t$. We also perturb with the correct orientation for a maximum principle to hold, that is, in directions out of the strip on the boundary. This is related to [Sei08, §(8k), Equation (8.21)] for moving boundary conditions as in continuation maps; in cases such as for a differential on $CF(K_{j},K_{j})$, one can perturb in the $t$-direction only. That is, one can use $\xi^{\#}\otimes dt$ near 0 so the perturbation is 0 on $\partial_{s}$ tangent to the boundary and $\xi^{\#}$ on $\partial_{t}$ normal to the boundary.
Analogous to [AA24, Equation (3.44)], define
Definition 3.9.
The continuation maps $F_{L_{0}^{k},L_{1}^{j}}:\mathcal{O}^{*}(L_{0}^{k},L_{1}^{j})\to\mathcal{O}^{*}(L_{0}^{k+1},L_{1}^{j+1})$ are
(3.22) $F_{L_{0}^{k},L_{1}^{j}}(p)=\sum_{q\in L_{0}^{k+1}\cap L_{1}^{j+1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in H_{2}(Y^{0},\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{0}^{k})\bigcup\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{1}^{j}))\\ \text{ind}(u)=0\end{subarray}}\#\mathcal{M}(p,q;[u],J)e^{-2\pi\int_{\beta}\omega}\cdot q$
(3.23) $\mathcal{M}(p,q;[u],J)$ $=\{u:(\mathbb{R}\times[0,1],\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\})\to(Y^{0},\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{0}^{k})\bigcup\cup_{s\in\mathbb{R}}\Lambda_{s}(L_{1}^{j})\mid$
$(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}=0,u(s,0)\subset\Lambda_{s}(L_{0}^{k}),u(s,1)\subset\Lambda_{s}(L_{1}^{j}),$
$\lim_{s\to\infty}u(s,t)=p,\lim_{s\to-\infty}u(s,t)=q\}$
where recall $\xi^{t}$ is the Hamiltonian vector field on the base dual to $dH$. Since $\rho^{t}$ is the flow of the horizontal lift $(\xi^{t})^{\#}$, then $(\xi^{\tau})^{\#}$ is a section of Hom($\Sigma,TY^{0})$ which at a point $z_{0}\in\Sigma$ is $(\xi^{\tau(z_{0})})^{\#}=\frac{d}{dt}\rho^{t}\mid_{t=\tau(z_{0})}$.
Theorem 3.10.
The moduli space $\mathcal{M}(p,q;[u],J)$ is a (1) smooth manifold, (2) compact, and (3) 0-dimensional, so it is a finite set of points which can be counted.
Proof.
These arguments are standard in the literature ([AD14, §8] for moduli spaces of cylinders between Reeb orbits, [MS12, §6.7] for domain dependent $J$, and [Sei08, §(9k), Equation (9.27)], [AS10, Lemma 8.5], [Sei12, §4] for discs). We adapt them to our setting. For 3), the moduli space is 0-dimensional because we count index 0 discs.
(1) The moduli space is the vanishing of the differential operator $u\mapsto(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$, thought of as a section of a Banach bundle, so by the Sard-Smale theorem it suffices to show the derivative (also known as the linearization) of the differential operator at each map $u$ in the moduli space, is surjective. (A finite-dimensional example of this process is the vanishing of $f(x,y,z)=x^{2}+y^{2}+z^{2}-a:\mathbb{R}^{3}\to\mathbb{R}$, which cuts out a smooth manifold for $a\neq 0$, a sphere, because its linearization $df=[2x\;2y\;2z]\neq\underline{0}$ at each such $(x,y,z)\in f^{-1}(0)$.)
We have two choices of datum, one for $\tau$ and one for $J$. The linearization of $(du+(\xi^{\tau})^{\#}\otimes d\tau)^{0,1}$ is
$(d\mathcal{F})_{u}(\dot{u},\dot{J},\dot{\tau})=D_{u}\dot{u}+\frac{1}{2}\dot{J}\circ(du+(\xi^{\tau})^{\#}\otimes\tau)\circ j+\dot{\tau}^{0,1}$
where the first two terms are justified in [Can20, Equation (4.15)] and the third term is from an infinitesimal perturbation of $\tau$.
Remark 3.11.
In the notation of [AD14, §8.4], their perturbation is $JX(s,t)$ so they linearize
$\mathcal{F}(u+Y)=\frac{\partial(u+Y)}{\partial s}+J(u+Y)\frac{\partial(u+Y)}{\partial t}-J(u+Y)X_{t}(u+Y)$
to the linear components
(3.24) $\frac{\partial Y}{\partial s}+(dJ)_{u}(Y)\frac{\partial u}{\partial t}+J(u)\frac{\partial Y}{\partial t}-(dJ)_{u}(Y)X_{t}-J(u)(dX_{t})_{u}(Y).$
In this paper, the perturbation term comes from $\tau(s,t)$ so instead of the last two terms in Equation (3.24) we have a single $\dot{\tau}^{0,1}$ term, what Seidel refers to as $(\delta Y)^{0,1}$ in [Sei08, Equation (9.26)], see also [AS10, Equation (8.10)] for perturbation term $X\otimes\gamma$ and linearization with respect to $u$ and $J$.
[…]
Recall that an injective point $z\in\mathbb{D}$ is one such that $d_{z}u\neq 0$ and $u^{-1}(u(z))=\{z\}$. We make use of the perturbation to ensure that discs are somewhere injective. Decomposing discs with Lagrangian boundary condition to somewhere injective discs has been studied in [KO00, Laz00, Laz11]. Here we show discs are already somewhere injective.
Lemma 3.12.
The set of injective points for a curve $u$ counted in Equation (3.23) is dense in $\mathbb{D}$.
Proof.
Discs aren’t multiply covered because $\tilde{u}:=W_{0}\circ u:\mathbb{R}\times[0,1]\to\mathbb{C}$ is an unramified holomorphic map of Riemann surfaces. The two boundary components map to different moving Lagrangians which prevents ramification points on the interior. A ramification point on the boundary means $d\tilde{u}$ collapses $\partial_{s}$ and $\partial_{t}=j\partial_{s}$ together but pseudo-holomorphicity implies one maps to the tangent space $TW_{0}(L)$ and the other to $J(TW_{0}(L))$. For a totally real submanifold, the only vector in their intersection $TW_{0}(L)\cap J(TW_{0}(L))=\{0\}$. Thus there couldn’t have been a ramification point on the boundary. So discs are not multiply covered, and now we justify why the set of injective points is dense.
To show non-injective points are isolated, first select $u:\Sigma\to Y^{0}$ a curve in the moduli space in Equation (3.23). Suppose by contradiction that we can pick a $z_{0}\in\Sigma$ such that there exists $z_{0}\neq z_{0}^{\prime}$ in $\Sigma$ and disjoint open neighborhoods of each, call them $U,U^{\prime}$ so that there is a biholomorphism $\varphi:U\to U^{\prime}$ where $u(z)=u\circ\phi(z)$ for all $z\in U$ but $z\neq\varphi(z)\in U^{\prime}$. Since $u|_{U}=u\circ\phi$ (and so $du|_{U}=du\circ d\phi$) agree on the neighborhoods $U,U^{\prime}$, but there exist vector $V$ and point $z$ where $(\xi^{\tau(z)})^{\#}d\tau_{z}(V)\neq(\xi^{\tau(\varphi(z))})^{\#}d\tau_{\varphi(z)}(d\varphi(V))$, then $u$ couldn’t have solved the Floer equation in Equation (3.23). The reason we introduced $\tau$ is to pick up the fact that $z\neq\phi(z)$ differ in the domain, even though from the perspective of the image they are the same. Note that $\tau$ cannot be constant by the definition of $\tilde{\chi}$ in Definition 3.8 and the definition of $\tau$ on the boundary by Equation (3.21). Thus non-injective points are isolated and injective points are dense. ∎
2) It remains to show the moduli space is compact. Since the moduli space of discs is a metric space, it suffices to show it is sequentially compact. Thus we need to show that every sequence has a convergent subsequence whose limit is in the space. This is done using Gromov compactness. To satisfy the assumptions of, and then apply, Gromov compactness, we split into two parts.
  1. (a)
    We need to exclude discs escaping to infinity. Then Gromov compactness applies and states there are only certain types of bubbling that could be in the limit.
  2. (b)
    We exclude those possible degenerations of curves.
2a) To exclude discs escaping to infinity, we need a maximum principle. Consider the annulus $A$ from Equation (3.4). Outside of a larger annulus, projected Lagrangian ends wrap according to the Hamiltonian in Equation (3.3).
Lemma 3.13.
For $u\in\mathcal{M}(p,q;[u],J)$, the image of the disc $u$ is contained in a compact subset of $Y^{0}$.
We adapt the methodology of [AA24, Proposition 3.10].
Proof.
First, let’s define the region we expect the curve to stay in within the base, because fibers are already compact. Enlarge $A$ to a larger annulus $\tilde{A}:=\{r_{1}\leq|w|\leq r_{2}\}\ni W_{0}(p),W_{0}(q)$ containing the points the projected strip limits to, and also $(R_{1},R_{4})\subset(r_{1},r_{2})$ so outside of $\tilde{A}$ projected Lagrangians have constant angle and isotopies only rotate them by Equation (3.3). We’d like to show $u(\mathbb{R}\times[0,1])$ is contained in the compact subset $\Omega:=Y^{0}|_{W_{0}^{-1}(\tilde{A})}$. We will do this by showing that maximums of $|W_{0}\circ u|$ and $|\frac{1}{W_{0}}\circ u|$, if they exist, must be attained on the boundary. Then we show that they cannot be attained on the boundary $\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\}$ outside $\Omega$ and $1/\Omega$ respectively. Therefore they must be attained in $\Omega$ or a maximum is approached at the limits to $p$ and $q$, both of which are contained in $\Omega$.
Recall that away from the singular fiber $J|_{Y^{0}\backslash\Omega}=J_{0}$ is multiplication by $i$ in the affine toric coordinates $t_{1},t_{2},t_{3}$ which are related to $y_{1},y_{2},y_{3}$ of Remark 1.4 by the process described in [ACL21, Equation (72)]. In particular, since $v_{0}=t_{3}$ we have $W_{0}=T^{\epsilon}(t_{3}-1)$ and so $W_{0}|_{\Omega}$ is $(J_{0},i)$-holomorphic because $dW_{0}\circ J_{0}=T^{\epsilon}dt_{3}\circ i=i\circ d(T^{\epsilon}(t_{3}-1))=i\circ dW_{0}$. Similarly $\frac{1}{W_{0}}:\mathbb{R}\times[0,1]\to\mathbb{C}^{*}$ is $(J_{0},i)$ holomorphic by composing with the $i$-holomorphic map $w\mapsto 1/w$.
We project $u$ to the base and absorb the isotopy $\xi^{\tau}$ so it satisfies the elliptic Cauchy-Riemann operator. Let $\tilde{\rho}^{t}$ be the flow of $\xi^{t}$, which is equal pointwise to $W_{0}\circ\rho^{t}\circ W_{0}^{-1}$ for any choice of $W_{0}^{-1}$. Then $\tilde{u}(s,t):=\tilde{\rho}^{\tau(s,t)}\circ W_{0}\circ u(s,t)$ is a $(j,(\rho^{\tau})_{*}j)$-holomorphic curve satisfying
(3.25) $(d\tilde{u})_{(j,(\rho^{\tau})_{*}j)}^{0,1}=\partial_{s}\tilde{u}+(\rho^{\tau})_{*}j\partial_{t}\tilde{u}=0$
where $j=i$ is the standard complex structure. Thus $\tilde{u}:\mathbb{R}\times[0,1]\to\mathbb{C}^{*}$ satisfies the maximum principle for holomorphic functions between open subsets of $\mathbb{C}$. If it has a maximum it must be on the boundary $\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\}$.
Let $j_{1}:=j$ and $j_{2}:=(\rho^{\tau})_{*}j$ so $\tilde{u}$ is $(j_{1},j_{2})$-holomorphic on domains in $\mathbb{C}$, thus in particular it is conformal for the metrics $g_{1},g_{2}$ compatible with $\omega_{\mathbb{C}}$ defined by $g_{k}(-,-)=\omega_{\mathbb{C}}(-,j_{k}(-))$ for $k=1,2$. Let $\tilde{u}:=r(s,t)e^{i\theta(s,t)}$ so $r(s,t)=r\circ\tilde{u}$ and $\theta(s,t)=\theta\circ\tilde{u}$. Then by expressing $d\tilde{u}$ in real coordinates $\tilde{u}=:\tilde{u}_{x}+i\tilde{u}_{y}$ and converting to polar coordinates, using Equation (3.25) we have
(3.26) $d\tilde{u}(\partial_{s})=-j_{2}\circ d\tilde{u}\circ j_{1}(\partial_{s})\implies\partial_{s}r(s,t)+ir(s,t)\partial_{s}\theta(s,t)=-j_{2}\partial_{t}r(s,t)-j_{2}ir(s,t)\partial_{t}\theta(s,t).$
We need to check that the disc boundary doesn’t stretch out infinitely along the Lagrangians under $\tilde{u}$. We start by showing boundedness away from infinity. Suppose there’s a maximum $r(s_{0},t_{0})$ on the boundary and outside of $\tilde{A}$, at some $(s_{0},t_{0})\in W_{0}(Y^{0})\backslash\{|w|\leq r_{2}\}$ where $t_{0}$ is 0 or 1. Then $\partial_{s_{0}}(r\circ\tilde{u})=0$ there. By the Hopf lemma, at a maximum, the derivative in the other direction $\partial_{t}$ moving outside the disc, must be positive. Because $e^{\pi z}$ is a biholomorphism from $\mathbb{R}\times(0,1)$ to the interior of the upper-half plane which extends to mapping $\mathbb{R}\times\{0\}\to\mathbb{R}_{+}$ and $\mathbb{R}\times\{1\}\to\mathbb{R}_{-}$, we can take coordinates on the strip to be parametrized by points on the upper-half plane (UHP) and then $\partial_{t^{UHP}}$ points outwards of the upper-half plane where $s^{UHP}+it^{UHP}$ are the UHP coordinates. The derivative of $r$ on $\tilde{u}$ must point outwards at $s_{0}^{UHP}+it_{0}^{UHP}:=e^{\pi(s_{0}+it_{0})}$ in the direction of $\partial_{t^{UHP}}$ so
$\frac{\partial(r\circ\tilde{u})}{\partial t^{UHP}}\mid_{(s_{0},t_{0})}<0.$
Therefore since $(\partial_{s},\partial_{t})$ is a positively-oriented $g_{1}$-orthogonal frame and $\tilde{u}$ is conformal then $(d\tilde{u}(\partial_{s}),d\tilde{u}(\partial_{t}))$ is a positively-oriented $g_{2}$-orthogonal frame and therefore
$\frac{\partial(\theta\circ\tilde{u})}{\partial s^{UHP}}\mid_{(s_{0},t_{0})}>0.$
In words, the image $\tilde{u}$ moves counterclockwise in the base $\mathbb{C}^{*}$ as we move in the direction of increasing $s$ on the boundary of the upper-half plane in the domain. From the definition of $\tilde{u}$, this movement has boundary on a fixed Lagrangian, either $W_{0}(L_{0}^{k})$ or $W_{0}(L_{1}^{j})$ depending on which connected component $(s_{0},t_{0})$ is on. This Lagrangian has ends which have strictly increasing radial coordinate outside the disc $\{|w|\leq r_{2}\}$ by Definitions 3.4 and 3.6. The tangent to the curve on the Lagrangian then has a positive radial component in the $s$ direction. Now we compare the two directions. Our analysis above showed that the curve $\tilde{u}$ has a positive angular component and zero radial component in the $s$ direction at $(s_{0},t_{0})$ which means it is rotated 90 degrees counterclockwise from the radial direction, and hence also counterclockwise from the Lagrangian boundary tangent vector at the same point. But, the isotopy from $\tau$ moves the Lagrangian clockwise. Since the two are going in opposite directions, this is a contradiction, and the curve $\tilde{u}$ could not achieve its maximum on the boundary outside $\tilde{A}$. For boundedness away from 0, we apply the above argument to $1/W_{0}$ on the inverted annulus $1/\tilde{A}$ and use Equation (3.10). ∎
2b) Now we exclude the possible types of bubbling described in Gromov compactness. We have the assumptions needed for Gromov compactness to hold by [AA24, Proposition 3.13] where their $X_{H}$ and $H$ terms are 0 in our setting since we don’t have wrapping in the fibers. The assumptions of their lemma hold because we proved in Lemma 3.13 above that $[u]$ is contained in a compact subset which we labeled $\Omega$ to match their notation.
Possible degenerations for discs are sphere bubbles (energy concentrates at an interior point), disc bubbles (energy concentrates at a boundary point), or strip-breaking (energy concentrates at a limiting intersection point). We can exclude sphere bubbles similar to [Can20, Corollary 4.53], because discs are somewhere injective by the paragraph above Lemma 3.12. Since Lagrangians $L$ are homeomorphic to $T^{2}$ and $\pi_{2}(T^{2})=0$, a disc bubble $u$ would have $u]\in\pi_{2}(T^{2},\rho^{t}(L))=0$ so must be constant. Therefore there can be no disc bubbles. Strip-breaking, from reparametrization by translation, is excluded because the prescribed moving Lagrangian boundary condition prevents translation as an automorphism; the continuation maps count index 0 strips which limit to fixed Lagrangian boundaries. So the moduli space cannot limit to any configurations not already in it and therefore is compact.
This concludes the proof of Theorem 3.10 that the moduli space $\mathcal{M}(p,q;[u],J)$ can be counted.
To justify the title of this subsection that continuation maps are inverses, we have the following lemma from [AA24], on cohomology. Composition in $\mu^{2}$ goes from right to left.
Lemma 3.14 ([AA24, Proof of Lemma 3.23]).
For $p_{k}\in HF^{*}(L_{0}^{k},L_{1}^{j})$ and $p_{k+1}\in HF^{*}(L_{0}^{k+1},L_{1}^{j})$
(3.27) $\mu^{2}(e_{L_{1}^{j}},p_{k+1})$ $=F_{L_{0}^{k},L_{1}^{j}}(\mu^{2}(p_{k+1},e_{L_{0}^{k}}))$
$\mu^{2}(F_{L_{0}^{k},L_{1}^{j}}(p_{k}),e_{L_{0}^{k}})$ $=\mu^{2}(e_{L_{1}^{j}},p_{k}).$
Taking them to represent elements in the quotients $\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k},L_{1}^{j})=\coprod_{j}HF^{*}(L_{0}^{k},L_{1}^{j})/\sim$ and $\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k+1},L_{1}^{j})=\coprod_{j}HF^{*}(L_{0}^{k+1},L_{1}^{j})/\sim$ respectively, the induced continuation maps on cohomology $F_{L_{0}^{k},L_{1}^{j}}:HF(L_{0}^{k},L_{1}^{j})\to HF(L_{0}^{k+1},L_{1}^{j+1})$ form an inverse to
$\mu^{2}(-,e_{L_{0}}^{k}):\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k+1},L_{1}^{j})\to\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k},L_{1}^{j}).$
The proof uses [AA24, Lemma 3.19] that Equation (3.27) holds up to homotopy on the chain-level by constructing parametrized moduli spaces [Sei08, p. 241], [Can20, Lemma 5.1] varying choices in 1 dimensional families and taking the boundary. It still applies in our setting because we have a maximum principle.
3.2.5. Morphisms in the derived category
We would like to glue together equivalent objects. A motivating example [Wei94, §1.5] is that of making a simplicial complex $X$ contractible by gluing it to a point. Taking its cone $CX$ then $C_{j}(CK)=C_{j-1}(K)\oplus C_{j}(K),\;j\geq 1$ from chains fully in the base versus those through the cone vertex $s$. Checking the differential and quotienting by $C_{0}$ of the cone point, this is just $C_{j}(CK)/C_{j}(\text{pt})=Cone_{j}(\mathbf{1}_{*}:C_{\bullet}(K)\to C_{\bullet}(K))$. So the notion of topological and algebraic cone coincide for topological spaces.
In practice when computing morphisms of Lagrangians in a Landau-Ginzburg model, the informal way of doing so is to “push the ends of the input Lagrangian, when projected to the base, to lie above the ends of the output Lagrangian.” In other words, if $L_{0}^{+}$ is $L_{0}$ with the ends pushed up, we would like a morphism $L_{0}\to L_{1}$ to be a “roof” $L_{0}^{+}$$L_{0}$$L_{1}$$\scriptstyle{e}$$\scriptstyle{p}$ where $e\in CF(L_{0}^{+},L_{0})$ is a quasi-unit and $p\in L_{0}^{+}\cap L_{1}$ is an intersection point or morphism of $CF(L_{0}^{+},L_{1})$. This roof is the standard construction of a left fraction in [Wei94, Prototype 10.3.5], which is a morphism in a localized category. To compose roofs would involve a pullback along $L_{0}^{+}\rightarrow L_{1}\leftarrow L_{1}^{+}$ [Wei94, Definition 10.3.4, Ore condition]. However, it is not clear what this pullback would be in practice.
We instead take a homotopy colimit over Yoneda modules $\mathcal{Y}_{L^{k}}=\mathcal{O}(-,L^{k})$ so we are in the abelian category of $\mathcal{O}$-modules. We follow [AA24, §3.5], where they use standard constructions similar to [AS10, §(3g)], [GPS20, §3.1.3]. These refer to [LO06] for localization of $A_{\infty}$-categories, a process that generalizes localization of categories [Wei94, §10.3], [KS06, §10.2], [Nee01, §1], [Rie14, Example 11.5.11], [BN93].
Definition 3.15 (Definition of morphisms).
Define
$\mathcal{Y}_{L^{\infty}}:=\text{hocolim}_{\mathcal{O}}\mathcal{Y}_{L^{k}}$
over the directed category $\mathcal{O}$ and morphisms
(3.28) $\mathcal{W}(L_{0},L_{1}):=\text{Hom}_{\mathcal{O}}(\mathcal{Y}_{L_{0}^{\infty}},\mathcal{Y}_{L_{1}^{\infty}}).$
So instead of Floer theory with $L$, we use $\mathcal{Y}_{L^{\infty}}$. This makes sense as the right object to use because we have a sequential diagram
$\ldots\mathcal{Y}_{L^{-1}}\to\mathcal{Y}_{L^{0}}\to\mathcal{Y}_{L^{1}}\to\ldots$
with connecting maps given by multiplication of quasi-units $\mu^{2}(e_{L^{k}},-):\mathcal{Y}_{L^{k}}\to\mathcal{Y}_{L^{k+1}}$. The homotopy colimit identifies objects that have a homotopy between them, preserving weak equivalences which are when a map of spaces induces an isomorphism on homotopy groups (so a particular inverse isn’t needed, as for a strong equivalence). This is similar to how a colimit glues along connecting maps, but a limit such as the pullback of $0\hookrightarrow[0,1]\hookleftarrow 1$ is empty even though $0,1,[0,1]$ are all homotopy equivalent, while the homotopy limit would be $\{\gamma:[0,1]\to[0,1]\mid\gamma(0)=0,\gamma(1)=1\}$, [Rie14, Exercise 6.5.4]. In the case of Lagrangians in the Fukaya category, that homotopy can be realized geometrically as an isotopy from one Lagrangian to another that more fully wraps. Furthermore, in HMS we take the split-closed derived (also called idempotent completion) of the Fukaya category, [Nee01, Proposition 1.6.8].
Now we justify why this definition for morphisms means we can push the ends of the input Lagrangian up. The following isomorphism implies that after taking cohomology of the chain level construction of $\mathcal{W}(L_{0},L_{1})$, we can take a sufficiently large wrapping of $k$ on the input Lagrangian in usual Floer theory.
Claim 3.16 ([AA24, Corollary 3.24]).
For each integer $k\in\mathbb{Z}$ and each pair $L_{0}^{k}$ and $L_{1}$ of objects of $\mathcal{O}$, there is a natural isomorphism
(3.29) $\text{colim}_{j\to\infty}HF^{*}(L_{0}^{k},L_{1}^{j})=\coprod_{j\in\mathbb{N}}HF^{*}(L_{0}^{k},L_{1}^{j})/(p\sim\mu^{2}(e_{L_{1}^{j}},p))\to H\mathcal{W}(L_{0},L_{1}).$
This holds because the following is a model for the homotopy colimit.
Claim 3.17 ([AA24, Equation (3.48)]).
(3.30) $\mathcal{Y}_{L^{\infty}}=\text{Cone}\left(\bigoplus_{k=0}^{\infty}\mathcal{Y}_{L^{k}}\xrightarrow[]{\oplus_{k=0}^{\infty}\mathbf{1}_{\mathcal{Y}_{L^{k}}}\oplus-\mu^{2}(e_{L^{k}},-)}\bigoplus_{k=0}^{\infty}\mathcal{Y}_{L^{k}}\right)$
This is a standard model from topology called the mapping telescope, adapted to abelian categories as described in [AS10, §(3g)], [BN93]. One can use it as the definition as in [Nee01, Definition 1.6.4]. More recent homotopical algebra developments, for example [Rie14, Example 11.5.11], explain why this agrees with the general definition of homotopy colimit as the left derived functor of the colimit constructed as the geometric realization of the cobar construction [Rie14, Corollary 5.1.3]. (A motivating example is the double mapping cylinder [Rie14, Example 6.4.5].)
To justify Equation (3.29), which is how we will compute morphisms in the next chapter, we see that it is a corollary of the following two results:
Claim 3.18 (Result 1, [AA24, Lemma 3.22]).
$\text{holim}_{k\to\infty}\text{hocolim}_{j\to\infty}\mathcal{O}(L_{0}^{k},L_{1}^{j})\to\mathcal{W}(L_{0},L_{1})$
is a quasi-isomorphism.
The proof starts with the mapping telescope model, and then relies on homological algebra properties which still apply in our setting. This leads to the following
Lemma 3.19 (Result 2, [AA24, Lemma 3.23]).
For all $L_{0},L_{1}$, and $k$, the map
$\operatorname{Hom}_{\mathcal{O}}(\mathcal{Y}_{L_{0}^{k+1}},\mathcal{Y}_{L_{1}^{\infty}})\to\operatorname{Hom}_{\mathcal{O}}(\mathcal{Y}_{L_{0}^{k}},\mathcal{Y}_{L_{1}^{\infty}})$
induced by multiplication by quasi-units is a quasi-isomorphism.
The proof follows again from homological algebra, the definition of $\mathcal{O}$, and Equation (3.27). These all still apply in our setting. Since the bonding maps in the holim of Result 1 are isomorphisms on cohomology by Result 2, taking a projection we find that for any $k$, Equation (3.29) holds.
3.2.6. Calculations
Here we take $j>i+1$. The results can be extended to all $i,j$ by Serre duality on the A- and B-side [ACLL23a, Lemma 4.7].
Lemma 3.20 (Morphisms on $\mathcal{F}_{1}$).
The Floer theory on the U-shaped Lagrangian objects of $\mathcal{F}_{1}$ is the same as it was for the same U-shaped Lagrangians of $(Y,-v_{0})$ ([Can20, Lemma 5.14]),
$HF^{*}(L_{i},L_{j})=\mathbb{C}^{(i-j)^{2}}/C\partial(\mathbb{C}^{(i+1-j)^{2}})$
where $\partial$ is the differential, $C$ is a scalar function of $x$ defined by an open Gromov-Witten invariant from [KL15] and $\partial$ as a linear map is proportional to the linear map $\vartheta\otimes$- on $\operatorname{Ext}(\mathcal{L}^{i+1},\mathcal{L}^{j})$, the defining multi-theta function of $\Sigma_{2}$.
Proof.
Note that U-shaped Lagrangians go below the puncture, and wrap for large $r(t)$, so they do not interact with the puncture. The moduli space in Equation (3.12) admits a regular $J$ because there exist a dense set of such $J$ following the roadmap from [MS12] outlined in [Can20, §4.5, Lemma 4.47], and restricting the $J$ there to $Y^{0}$ here. Thus the moduli space counted in [Can20]
(3.31) $\mathcal{M}(p,q;\beta,J):=\left\{u:\mathbb{R}\times[0,1]\to Y\middle|\begin{array}[]{l}[u]=\beta,\ \overline{\partial}_{J}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{i},\ u(\mathbb{R},1)\subset L_{j},\\ u(-\infty)=q,\ u(+\infty)=p\end{array}\right\}/(u\sim u_{a})$
is isomorphic by a 1-1 bijection with the moduli space counted here (restricting $Y$ and $J$ to $Y_{0}$)
(3.32) $\mathcal{M}(p,q;\beta,J|_{Y^{0}}):=\left\{u:\mathbb{R}\times[0,1]\to Y^{0}\middle|\begin{array}[]{l}[u]=\beta,\ \overline{\partial}_{J|_{Y^{0}}}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{i},\ u(\mathbb{R},1)\subset L_{j},\\ u(-\infty)=q,\ u(+\infty)=p\end{array}\right\}/(u\sim u_{a}).$
Remark 3.21.
Because we are in the non-exact setting, changes in area affect calculations. In particular, homotoping $\gamma_{L_{j}}$ and $\delta_{K_{j}}$ (staying away from the critical values) to different choices then isotopes the Lagrangians above them and affects disc area calculations in the Fukaya category according to the formula [ACLL23a, Proposition 6.10]. However, in a symplectic Landau-Ginzburg model, the calculations split into fiber and base contributions because the Lagrangian isotopies (coming from parallel transport over homotopies) are exact [ACLL23b, Lemma 2.1], and different choices are therefore absorbed into $C$ by [ACLL23b, Corollary 2.4].
For morphisms between $\mathcal{F}_{1},\mathcal{F}_{2}$, suppose $\gamma_{L_{i}}\cap\delta_{K_{j}}$ intersect in two points as depicted in Figure 3 which are
$w_{1}:=\gamma_{L_{i}}(t_{1})=\delta_{K_{j}}(s_{1}),\quad w_{2}:=\gamma_{L_{i}}(t_{2})=\delta_{K_{j}}(s_{2}).$
There are bigons bounded by Lagrangians $L_{i}\cup K_{j}$ covering the bigon in the base between these two paths and intersection points. Hamiltonian-perturb that bigon to a circle. That is, let
$\zeta:S^{1}\to\mathbb{C}^{*}$
denote the projection of a portion of Hamiltonian-equivalent Lagrangians $\phi_{H}(L_{i}\cup K_{j})$ which project to a circle centered at $T^{\epsilon}$ and let $\zeta(1)$ be the image of $w_{1}$ under the Hamiltonian diffeomorphism composed with $W_{0}$. Then define (where subscripts $Y_{w}:=Y|_{W_{0}^{-1}(w)}$ denote the fiber over that point)
$\Phi^{s}:Y^{0}_{w_{1}}\to Y^{0}_{\zeta(1)}\to Y^{0}_{\zeta(e^{2\pi is})}\to Y^{0}_{\phi_{H}^{-1}(\zeta(e^{2\pi is}))}$
to be the composition of the Hamiltonian isotopy, parallel transport counterclockwise around the circle an angle of $s$ from the first intersection point at $w_{1}$, and then reversing the Hamiltonian isotopy back to the original Lagrangians. Lastly, let
$\ell_{i}^{\prime}:=\Phi_{\gamma_{L_{i}}(t_{1})}(\ell_{i}),\qquad\ell_{j}^{\prime\prime}:=\Phi_{\delta_{K_{j}}(s_{1})}(\ell_{j}),\qquad\Phi:=\Phi^{1}$
so $\Phi=\Phi^{1}$ denotes the monodromy of $W_{0}$ around the singular fiber. With this set-up, we can now examine intersections between $L_{i}$ and $K_{j}$.
Lemma 3.22.
$CF(L_{i},K_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$.
Proof.
As depicted in Figure 3,
(3.33) $CF_{Y^{0}}(L_{i},K_{j})$ $=HF_{Y^{0}_{w_{1}}}(\ell^{\prime}_{i},\ell^{\prime\prime}_{j})[1]\oplus HF_{Y^{0}_{w_{2}}}(\Phi^{1-s}(\ell^{\prime}_{i}),\Phi^{-s}(\ell^{\prime\prime}_{j}))$
$=HF_{T^{4}}(\ell_{i},\ell_{j})[1]\oplus HF_{T^{4}}(\Phi^{1}(\ell_{i}),\ell_{j})$
$=HF_{T^{4}}(\ell_{i},\ell_{j})[1]\oplus HF_{T^{4}}(\ell_{i+1},\ell_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$
where we use $HF$ on the righthand-side because in a fiber $\partial=0$ as there are no bigons between linear Lagrangians in a 4-torus. We start with $\ell^{\prime}_{i},\ell^{\prime\prime}_{j}$ in the fiber over the left intersection point, and parallel transport around the loop around the origin with respect to that starting point. The second equality follows because $\Phi^{s}$ is a diffeomorphism for any $s$ so preserves intersection points, hence we can apply $\Phi^{s}$ to Lagrangians in the second factor of the $\oplus$ and then note that $\Phi^{1}(\ell_{i})$ is Hamiltonian isotopic to $\ell_{i+1}$ [Can20, Lemma 4.24] so the statement follows from the Hamiltonian invariance of $HF$ [Aur14, Theorem 1.5]. The gradings will be justified in Subsection 4.1.
S-S
0
$\infty$
$\times$
[…]
$L_{i}$
$\phi(L_{i})$
$\ell^{\prime}_{i}\cap\ell^{\prime\prime}_{j}$
$\Phi^{s}(\Phi^{(1-s)}(\ell^{\prime}_{i})\cap\Phi^{-s}(\ell^{\prime\prime}_{j}))$
$\Phi^{1}(\ell^{\prime}_{i})\cap\ell^{\prime\prime}_{j}$
(a) The base as a cylinder. The Lagrangian intersection points of $CF(L_{i},K_{j})$ are denoted in the two fibers over the base intersections, upon positively wrapping the ends of $L_{i}$. Here from the left intersection point to the right, we go counterclockwise $2\pi(1-s)$ around $W_{0}(\phi(L_{i}))$ and clockwise $2\pi s$ around from $W_{0}(K_{j})$. So up to diffeomorphism the differential counts bigons between two fibers, $\ell_{i}\cap\ell_{j}$ and $\Phi(\ell_{i})\cap\ell_{j}$ where $\Phi$ denotes monodromy around the singularity, and $\Phi(\ell_{i})\cong\ell_{i+1}$.
$\times$
0
[…]
$L_{i}$
$\phi(L_{i})$
$\mathbb{C}^{*}$
(b) The base as $\mathbb{C}^{*}$: projection of indicated Lagrangians under $W_{0}$
Figure 3. Projection of $K_{i}$, $L_{j}$, and the positively wrapped $\phi(L_{j})$ to the base of $W_{0}:Y^{0}\to\mathbb{C}$. There is a puncture at 0, and the singular fiber is at $T^{\epsilon}$, denoted by $\times$.
Now we compute the differential and pass to cohomology.
Lemma 3.23 ($\operatorname{Hom}(\mathcal{F}_{1},\mathcal{F}_{2})$).
Recall $CF(L_{i},K_{j})\cong CF(\ell_{i},\ell_{j})[1]\oplus CF(\ell_{i+1},\ell_{j})\cong\mathbb{C}^{(i-j)^{2}}\oplus\mathbb{C}^{(i+1-j)^{2}}$. The differential can be computed as in [Can20, Lemma 5.14] so Floer groups are again defined from
$CF(\ell_{i+1},\ell_{j})$ $\xrightarrow{\partial}CF(\ell_{i},\ell_{j})[1]\to HF(L_{i},K_{j})\to 0$ to be
$HF(L_{i},K_{j})=\mathbb{C}^{(i-j)^{2}}/C\partial(\mathbb{C}^{(i+1-j)^{2}}).$
Proof.
Since the moduli spaces are isomorphic, we obtain the same count as in Lemma 3.20 which was for two U-shaped curves. An illustration of how the two moduli spaces can be identified is given in Figure 4. ∎
$\times$
0
[…]
$\phi(L_{i})$
$\mathbb{C}^{*}$
$Y^{0}\hookrightarrow Y$
$\phi^{\prime}(K_{j})$
$\times$
$\mathbb{C}$
Figure 4. The differential calculation between one U-shape and one arc on $Y^{0}$ is equivalent to the differential calculation between two U-shapes on $Y$ by including $Y^{0}\hookrightarrow Y$, applying a Lagrangian isotopy $\phi^{\prime}$ on the arc $K_{j}$ by homotoping the curve in the base as shown in solid line, and then completing it to a U-shape as shown with the additional dotted line.
Lemma 3.24 ($\operatorname{Hom}(\mathcal{F}_{2},\mathcal{F}_{1})$).
Swapping the order of the objects, $HF(K_{j},L_{i})=0$.
Proof.
Morphisms are defined by positive isotopies of the input Lagrangian that wrap arbitrarily close to $\pi$ around $\infty$ in the $\mathbb{C}^{*}$ model of the base by Equation (3.29). Thus the end of the projected input Lagrangian will wrap past the ends of the projected output Lagrangian. In other words, we displace $K_{j}$ from $L_{i}$ in the base so that it is a ray that lies entirely above $L_{i}$ in the base. Since it is already entirely above $L_{i}$ to start, there are no intersection points in the base so there can be no intersection points in the total space. ∎
This exhibits the semi-orthogonality between $\mathcal{F}_{1},\mathcal{F}_{2}$ because there are morphisms in one direction and none in the other.
Lemma 3.25 (Morphisms on $\mathcal{F}_{2}$).
$HF(K_{i},K_{j})\cong\mathbb{C}^{(j-i)^{2}}[x].$
Proof.
Intersection points in the total space cover intersections in the base. By Equation (3.3), we only fully wrap at one end of the cylinder. Counting intersections and triangles in the base cylinder (a standard calculation [Abo10], [Aur14, §4.2]) we find that $HF(W_{0}(K_{i}),W_{0}(K_{j}))\cong\mathbb{C}[x]$ as algebras. In the torus fiber $CF(\ell_{i},\ell_{j})=HF(\ell_{i},\ell_{j})=\mathbb{C}^{(j-i)^{2}}$ ([Fuk02], [ACLL, §4.2]). Thus over each intersection point corresponding to $x^{k}$ in the base we have $\mathbb{C}^{(j-i)^{2}}$ intersection points in the fiber and
$HF(K_{i},K_{j})=CF(\ell_{i},\ell_{j})\otimes\mathbb{C}[x]\cong\mathbb{C}^{(j-i)^{2}}[x].$
4. Product structure
We compose from inside to out, meaning from right to left. Due to orthogonality, composition on the B-side is nonzero in two cases not covered above; either the first morphism is on the $H$ factor and then we compose with a morphism from $H$ to the $V\times\mathbb{C}$ factor, or the first morphism is from the $H$ factor to the $V\times\mathbb{C}$ factor and we compose with one on $V\times\mathbb{C}$. In other words,
$\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))\otimes\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)))$
$\hskip 72.26999pt\to\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
and
$\operatorname{Ext}(\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}}))\otimes\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}}))$
$\hskip 72.26999pt\to\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{k_{3}}\boxtimes\mathcal{O}_{\mathbb{C}})).$
These can be simplified using the isomorphism of Equation (2.3)
$\operatorname{Ext}(p^{*}i^{*}(\mathcal{L}^{k}\boxtimes\mathcal{O}_{\mathbb{C}}),j^{*}\pi^{*}(\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}})[-1])\cong\operatorname{Ext}(\mathcal{L}^{k}|_{H},\mathcal{L}^{j}|_{H})[-1]$
under which we restrict to $H\times\{0\}$. For example, sections of $\mathcal{O}_{\mathbb{C}}$ are polynomials, and under this isomorphism they are evaluated at 0. Sections of $\mathcal{L}^{j}$ for $j>0$ are multi-theta functions, infinite series, as in [ACLL, Equation (2.14)].
When composing morphisms between factors on the A-side, of the three Lagrangians, we see similar behavior; by the geometry of wrapping we can have either two which are arcs and one which is U-shaped, or two which are U-shaped and one which is an arc. For this we would like to use the result of [ACLL23b, Corollary 2.6], therefore we need to prove the assumption in the corollary, that given a pseudo-holomorphic curve $u$ counted in the composition map over a triangle away from the singularity, and given the map $u^{\prime}$ defined in [ACLL23a, Proposition 6.10] obtained by moving one Lagrangian under a Lagrangian isotopy so $\partial u^{\prime}$ is in a fiber, then $u^{\prime}]$ is relatively homologous to a pseudo-holomorphic triangle.
4.1. Grading
Let $L$ be $L_{j_{2}}$ (respectively $K_{j_{2}}$). For $\mathcal{D}(Y^{0},(L_{j_{1}},L,K_{j_{3}}),(p_{0},p_{1},p_{2}))$ the class of discs mapping to the Lagrangian boundary conditions (see [ACLL23a, Definition 5.1]), let
$\mathcal{M}(Y,(L_{j_{1}},L,K_{j_{3}}),(p_{0},p_{1},p_{2});\beta,J)\\ =\{u\in\mathcal{D}(Y^{0},(L_{j_{1}},L,K_{j_{3}}),(p_{0},p_{1},p_{2}))\mid[u]=\beta,\overline{\partial}_{J}(u)=0\}$
be the moduli space of $J$-holomorphic discs in the composition
(4.1) $HF(L,K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L)\to HF(L_{j_{1}},K_{j_{3}}[-1])$
depicted under projection of $W_{0}$ in Figure 7 when $L=L_{j_{2}}$ (respectively in Figure 9 when $L=K_{j_{2}}$).
The typical grading in a bigon in $\mathbb{R}^{2}$ has degree 0 on the right intersection point and degree 1 on the left intersection point, as in Figure 5 and [Aur14, Example 1.9]. When grading the intersection points in a Floer complex $CF(L_{1},L_{2})$ in the total space of a Landau-Ginzburg model, the fiber and base contributions add by [ACLL23b, Lemma 3.11 and Corollary 3.12]. The fiber case is covered in [ACLL23a, Example 4.6], where $\ell_{k}$ is linear of slope $k$ and the Lagrangian grading is $\frac{1}{\pi}\arctan(k)$. So here we focus on the base. We make a continuous choice of angle $\alpha_{j}(t)$ along the projected Lagrangian $W_{0}(L_{j})$ for a tangent vector to $T_{\gamma_{L_{j}(t)}}W_{0}(L_{j})$ relative to the positive horizontal line field, indicated in Figures 7 and 9 by arrows.
To find the degree of an intersection point in the base, we can break it up into three steps. Step 1 is to find $\alpha_{2}(t)-\alpha_{1}(t)$ at values of $t$ where the paths intersect (direction of the arrows on the base Lagrangians matter). Step 2: at that same value of $t$, we rotate the line $T_{\gamma_{L_{2}(t)}}W_{0}(L_{2})$ counterclockwise until it coincides with that of $T_{\gamma_{L_{1}(t)}}W_{0}(L_{1})$ (direction of the arrows on the base Lagrangians don’t matter). Since counterclockwise is always positive, we add that to $\alpha_{2}(t)-\alpha_{1}(t)$. This is referred to as the reverse canonical short path. Step 3: Twice this winding number is the degree, because rotating a line back to itself, through an angle $\pi$, should count as one rotation.
Remark 4.1.
Note that $(\exp(2\pi i\alpha_{j}(t)))^{2}=\alpha_{L}^{\text{Hor}}(p)$ of [ACLL23b, Lemma 3.11, Eq. (3.8)] where, for horizontal lift $(\dot{\gamma}_{L_{j}}(t))^{\#}$ and recalling $r(t)$ and $\theta(t)$ parametrizing $L_{j}$ as in Definition 3.4,
(4.2) $\alpha_{L}^{\text{Hor}}(p)$ $=\frac{(dW_{0}(\dot{\gamma}_{L_{j}}(t)^{\#}))^{2}}{|dW_{0}(\dot{\gamma}_{L_{j}}(t)^{\#})|^{2}}=\frac{(\dot{\gamma}_{L_{j}}(t))^{2}}{|\dot{\gamma}_{L_{j}}(t)|^{2}}=\frac{(\dot{r}(t)+r(t)2\pi\mathbf{i}\dot{\theta}(t))^{2}e^{2\pi\mathbf{i}\cdot 2\theta(t)}}{|\dot{r}(t)+r(t)2\pi\mathbf{i}\dot{\theta}(t)|^{2}}=e^{2\pi\mathbf{i}\cdot 2\alpha_{j}(t)}.$
For example for a circle $\dot{r}(t)=0$, $\dot{\theta}(t)=1$ and
$e^{2\pi\mathbf{i}\cdot 2\theta(t)}=e^{2\pi\mathbf{i}\cdot 2\alpha_{j}(t)}$
implies one choice for $\alpha_{j}(t)$ is
$\alpha_{j}(t)=\theta(t)+1/2.$
In contrast to the typical grading, here we want choices so that the three bigons involved in each differentials in the three $HF$ factors appearing in the composition in Equation (4.1) have degree 0 on the left intersection point and degree 1-1 on the right intersection point. This is done in order to mirror the shift in Ext groups on the B-side. That is, U-shapes in the base are all graded the same which means in the bigon, the choice of $\alpha_{j_{2}}(t)$ is shifted by $\pi$ from Figure 5, as in Figure 7. Replace the choice of $\alpha_{2}(t)$ in Figure 5 with $\alpha_{j_{2}}(t)=\alpha_{2}(t)-\pi$ for the definition of $L_{j_{2}}$. We do a similar adjustment for $K_{j_{3}}$ so that $\alpha_{j_{3}}(t)=-\pi$. Thus $K_{j_{3}}$ is replaced with $K_{j_{3}}[-1]$.
Labeling the base intersection points as $b_{0},\ldots,b_{5}$, we calculate the degrees $\deg(b_{j})$. Explanation is given in Figure 8 for $b_{0}$ and $b_{1}$. The other points are calculated similarly.
[…]
Figure 5.
$L_{j_{1}}$
$L_{j_{2}}$
$K_{j_{3}}$
$\deg(b_{1})=-1$
$\deg(b_{0})=0$
$\deg(b_{2})=0$
$\deg(b_{3})=-1$
$\deg(b_{4})=0$
$\deg(b_{5})=-1$
$\times$
$b_{0}$
$b_{1}$
$b_{2}$
$b_{3}$
$b_{4}$
$b_{5}$
Figure 6. Composition $HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1])$ depicted in the base in the cylinder model. The thicker arrow corresponds to the grading on the input projected Lagrangian of the Floer group.
Figure 7.
[…]
Figure 8. : Add $\frac{1}{2}$ because the oriented tangent vector rotates counterclockwise $\frac{\pi}{2}$ from $b_{0}$ to $b_{1}$
x
o
$\epsilon^{\prime}$
$\deg(c_{1})=0$
$\deg(c_{0})=0$
$\deg(c_{2})=0$
$\deg(c_{3})=-1$
$\deg(c_{4})=-1$
$c_{1}$
$c_{3}$
$c_{0}$
$c_{2}$
$c_{4}$
Figure 9. Composition $HF(K_{j_{2}}[-1],K_{j_{3}}[-1])\otimes HF(L_{j_{1}},K_{j_{2}}[-1])\to HF(L_{j_{1}},K_{j_{3}}[-1])$ depicted in the $\mathbb{C}^{*}$ base model, denote $D$ as the shaded region
Lemma 4.2.
In the moduli space defining
$HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1]),$
consider
(4.3) $u\in\mathcal{M}(Y,(L_{j_{1}},L_{j_{2}},\Phi_{H}(K_{j_{3}}[-1])),(p_{0},p_{1},p_{2});\beta,J)$
where $\Phi_{H}:Y^{0}\to Y^{0}$ is a Hamiltonian diffeomorphism to be defined in the proof and $W_{0}(\beta)$ covers the shaded region in Figure 7. Let $u^{\prime}$ be obtained by Lagrangian isotopy of $K_{j_{3}}$ so that all three intersection points pass through the same fiber, following [ACLL23b, Corollary 2.6]. Then $u^{\prime}]$ is relatively homologous to a pseudo-holomorphic triangle $u^{\prime\prime}$ in the fiber bounded by $\ell_{j_{1}}\cup\ell_{j_{2}}\cup\ell_{j_{3}}$.
The point is, we want to get a handle on $u$ to consider composition on the A-side, but there is ambiguity in how we draw the Lagrangians. With this Lemma and then subsequently [ACLL23b, Corollary 2.6], we can move everything to a fiber triangle $u^{\prime\prime}$, and then count the triangles by area from HMS on tori [ACLL, §4.3], [Fuk02].
Proof.
We construct a singular 3-chain, whose boundary is $u^{\prime}]-[u^{\prime\prime}]$ where $u^{\prime\prime}$ is in a fiber. The 3-chain is a continuous map from the unit cube to $Y^{0}$, and a cube can be triangulated into 3-simplices. For simplicity, denote $W_{0}$ as $v$.
Let $\delta(s,t):I\times I\to\mathbb{C}$ for $I=[0,1]$ parametrize the projection of the triangle $(v\circ u)(\mathbb{D})$ so that $\delta:(0,1)\times(0,1)\to(v\circ u)(\text{int}(\mathbb{D}))$ is a homeomorphism, $\delta(1,t),0\leq t\leq 1$ is the top side of the base triangle in Figure 7 before the isotopy, and the isotopy happens over the curve $\delta(s,t),0\leq s\leq 1$ for each $t$ (this is consistent with the use of $\delta$ in [ACLL23b, Corollary 2.6]). Then for $\Phi_{\gamma(0)\to\gamma(1)}$ denoting parallel transport along a path $\gamma$ between the fibers over $\gamma(0)\to\gamma(1)$ (as in [ACLL23b, Equation (1.3)]), define the singular 3-chain
(4.4) $Q:I\times I\times I$ $\to Y$
[…]
where
$D:=(v\circ u)(\mathbb{D})$
and $u_{D}:D\to Y^{0}$ is defined by
$u_{D}(\delta(s,t))=u(z(s,t))$
where $z(s,t)\in\mathbb{D}$ is uniquely defined by
$(v\circ u)(z(s,t))=\delta(s,t).$
because $u$ is 1-1. This is because we are working with a Landau-Ginzburg model [Sei08, Equation (17.2)] where curves in structure maps are sections of the superpotential $v$. In other words, $u_{D}:D\to Y^{0}$ is a $((v\circ u)_{*}j,J)$-holomorphic section of $v$ if $u$ was a $(j,J)$-holomorphic curve.
Then $Q$ is a continuous map by definition. The image of $Q$ is depicted in [ACLL23b, Figure 2], the prism over the middle image. Its boundary is (where paranthetical comments refer to [ACLL23b, Figure 2])
(4.5) $z=0,\quad 0\leq s,t\leq 1$ $\mapsto[u]\text{ (bottom triangle) }$
$z=1,\quad 0\leq s,t\leq 1$ $\mapsto[u^{\prime\prime}]\text{ (fiber triangle) }$
$s=0,\quad 0\leq z,t\leq 1$ $\mapsto[u^{\prime}]-[u]\text{ (top quadrilateral obtained from the Lagrangian isotopy) }$
$s=1,\quad 0\leq z,t\leq 1$ $\mapsto u(\delta(1,-))\text{ (constant at bottom vertex) }$
$t=0,1,\quad 0\leq s,z\leq 1$ $\mapsto K_{j_{2}}\cup K_{j_{3}}\text{ (front and back triangles) }$
where $u^{\prime\prime}$ is defined as follows. In the case of Figure 7, define the parameters along the boundary of the triangle in the base by
(4.6) $\gamma_{L_{j_{1}}}(t_{L_{j_{1}}}(s))$ $=\delta(s,1)$
$\gamma_{L_{j_{2}}}(t_{L_{j_{2}}}(s))$ $=\delta(s,0)$
$\gamma_{K_{j_{3}}}(t_{K_{j_{3}}}(t))$ $=\delta(0,t).$
This clarifies the relation between the homotopy $\delta(s,t)$ and the parameters defining the base curves for the Lagrangians. Then the $z=1$ face has boundary on one side on the Lagrangian in the fiber over S-S given by
(4.7) $\Phi_{\gamma_{L_{j_{1}}}(t_{L_{j_{1}}}(s))\to\gamma_{L_{j_{1}}}(0)}\circ\Phi_{\gamma_{L_{j_{1}}}(0)\to\gamma_{L_{j_{1}}}(t_{L_{j_{1}}}(s))}(\ell_{j_{1}})=\ell_{j_{1}}$
and another side in the same fiber in the Lagrangian
(4.8) $\Phi_{\gamma_{L_{j_{2}}}(t_{L_{j_{2}}}(s))\to\gamma_{L_{j_{2}}}(0)}\circ\Phi_{\gamma_{L_{j_{2}}}(0)\to\gamma_{L_{j_{2}}}(t_{L_{j_{2}}}(s))}(\ell_{j_{2}})=\ell_{j_{2}}.$
For the third side, we apply a Hamiltonian diffeomorphism $\Phi_{H}$ defined first on fibers over the initial third side
(4.9) $\Phi_{H}=\Phi_{\delta(1,t)\to\delta(0,t)}\circ\Phi_{\delta_{K_{j_{3}}}(0)\to\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}^{-1}:\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$ $\mapsto\bigcup_{0\leq t\leq 1}Y^{0}_{\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}$
$\Phi_{\delta_{K_{j_{3}}}(0)\to\delta_{K_{j_{3}}}(t_{K_{j_{3}}}(t))}(p)$ $\mapsto\Phi_{\delta(1,t)\to\delta(0,t)}(p).$
Because the composition of the two maps is parallel transport around a closed loop not enclosing the singular fiber by Equation (4.6), the composition is Hamiltonian (for the submanifold given by the fibers over the $\delta(0,t)$ part of the triangle), by [MS98, Theorem 6.21 (ii)]. This defines a Hamiltonian function $H$ on that submanifold, which we can then extend to the whole space smoothly. We call the resulting Hamiltonian diffeomorphism $\Phi_{H}$. The reason for applying this Hamiltonian diffeomorphism is so that under the $\delta$ isotopy, the third side is in the S-S fiber again in a linear Lagrangian
(4.10) $\Phi_{\delta(0,t)\to\delta(1,t)}\circ\Phi_{\delta(1,t)\to\delta(0,t)}(\ell_{j_{3}})=\ell_{j_{3}}.$
Looking at the universal cover of $(T^{4},\ell_{j_{1}}\cup\ell_{j_{2}}\cup\ell_{j_{3}})$, there is a unique flat holomorphic triangle in each homology class, which $\beta$ determines. Let $u^{\prime\prime}$ be that triangle.
Thus
$\partial Q=0=[u^{\prime\prime}]-[u^{\prime}]$
in $H_{2}(Y^{0},L_{j_{1}}\cup L_{j_{2}}\cup\psi(\Phi_{H}(K_{j_{3}})))$ where $\psi$ is the Lagrangian isotopy induced by $\delta$ and so
$u^{\prime}]=[u^{\prime\prime}]$
where $u^{\prime\prime}$ is in a $T^{4}$-fiber bounded by the three linear Lagrangians. ∎
The proof for 2 cotangent fibers and 1 U-shape as in Figure 9 is similar, where $\delta(s,t)$ is a homotopy moving the right side of the triangle, $L_{j_{1}}$ across to $\epsilon^{\prime}$ and $W_{0}(\beta)$ covers again the shaded region. Now $\Phi_{H}$ is chosen to have $L_{j_{1}}$ match the perturbed $K_{j_{2}}$ and $K_{j_{3}}$. Then in a fiber the triangle is bounded by $\Phi_{\delta_{K_{j}}(0)\to\epsilon^{\prime}}(\ell_{j_{1}}\cup\ell_{j_{2}}\cup\ell_{j_{3}})$ from going from S-S to $\epsilon^{\prime}$ along the path for $W_{0}(K_{j})$ in Definition 3.6. We can apply the symplectomorphism $\Phi_{\delta_{K_{j}}(0)\to\epsilon^{\prime}}^{-1}$ to obtain a bijection between those triangles and the linear ones, that preserves their weights.
Lemma 4.3.
The compositions with two U-shaped Lagrangians and one cotangent fiber Lagrangian
$HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1])$
and with two cotangent fiber Lagrangians and one U-shaped Lagrangian
$HF(K_{j_{2}}[-1],K_{j_{3}}[-1])\otimes HF(L_{j_{1}},K_{j_{2}}[-1])\to HF(L_{j_{1}},K_{j_{3}}[-1]),$
after rescaling, are products in the abelian variety fiber.
Proof.
One cotangent fibered and two U-shaped: each Floer cochain complex is a direct sum of fiber cochain complexes over two base as can be seen in Figure 7. For example
(4.11) $CF(L_{j_{2}},K_{j_{3}}[-1])=CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\oplus CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})$
in fibers over $b_{4}$ and $b_{5}$. The differential for each cochain complex goes from the intersection point on the right (degree 1-1) to that on the left (degree 0). For example, we have an exact sequence computing Floer cohomology from
(4.12) $0\to CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})[-1]\xrightarrow[]{\partial}CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\to HF(L_{j_{2}},K_{j_{3}}[-1])\to 0$
where $\partial$ counts images of bigon sections of $W_{0}$ over the bigon in the base between $b_{4}$ and $b_{5}$.
Lemma 4.2 showed that composition $M^{2}$ of the degree 0 components
(4.13) $M^{2}:CF^{0}(L_{j_{2}},K_{j_{3}}[-1])\otimes CF^{0}(L_{j_{1}},L_{j_{2}})$ $\to CF^{0}(L_{j_{1}},K_{j_{3}}[-1])$
$M^{2}:CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\otimes CF((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})$ $\to CF((L_{j_{1}})_{b_{2}},(K_{j_{3}}[-1])_{b_{2}})$
can be computed in the fiber with linear Lagrangians. For degree reasons, given that $M^{2}$ is of degree 0, there are only two other possibilities for the triangles:
(4.14) $M^{2}:CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})\otimes CF((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})\to CF((L_{j_{1}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})$
and
(4.15) $M^{2}:CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\otimes CF((L_{j_{1}})_{b_{1}},(L_{j_{2}})_{b_{1}})\to CF((L_{j_{1}})_{b_{3}},(K_{j_{3}}[-1])_{b_{3}}).$
These show that $M^{2}$ passes to cohomology, as follows. Recall that $M^{2}$ satisfies a Leibniz rule, in that modulo 2,
(4.16) $\partial\circ M^{2}(\cdot,\cdot)=M^{2}(\partial(\cdot),\cdot)+M^{2}(\cdot,\partial(\cdot)).$
Applying this to Equation (4.14), we find that for $c\in CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})$ and $c^{\prime}\in CF((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})$
(4.17) $\partial\circ M^{2}(c,c^{\prime})=M^{2}(\partial(c),c^{\prime})+\cancel{M^{2}(c,\partial(c^{\prime}))}$
where the second term is zero because $\partial(c^{\prime})=0$ since there $\partial$ increase degree by 1, $c^{\prime}$ is degree 0, and there are no intersection points of degree 1. By the analogous calculation for Equation (4.15), we find that $M^{2}$ passes to the quotient by the image of $\partial$, to cohomology:
(4.18) $HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1])$
and it can be computed with representatives in the degree 0 components, whose product can be reduced to products in the abelian variety fiber. A similar argument can be done in the second case.
To show that we can make the products equal that of the fiber products by rescaling the intersection points, first consider the first case. There, the Lagrangian $K_{j_{3}}$ is moved, which moves $b_{2}$ (output of $M^{2}$) and $b_{4}$ (an input of $M^{2}$). Pick intersection points over these base points, $p_{1}\in Y^{0}_{b_{4}}$ the fiber over $b_{4}$, and $q\in Y^{0}_{b_{2}}$. Then
$\rho(u)\rho(u^{\prime})^{-1}=\rho(u)\rho(u^{\prime\prime})^{-1}=e^{2\pi if(p_{1})}e^{-2\pi if(q)}$
by the last equation in the proof of [ACLL23b, Corollary 2.6] for a certain function $f$ defined there and depending on the Lagrangian isotopy. In other words, the original triangle and the fiber triangle differ an amount that only depends on the intersection points $p_{1},q$. Thus rescaling $p_{1}\mapsto e^{2\pi if(p_{1})}p_{1}$ and $q\mapsto e^{2\pi if(q)}q$, we find that for $p_{0}\in CF^{0}((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})$
(4.19) $p_{1}\cdot p_{0}=\sum_{q\in L_{j_{1}}\cap K_{j_{3}}}C_{p_{0}p_{1}q}q$
where $C_{p_{0}p_{1}q}$ is the count of triangles between linear Lagrangians in a fiber. The proof for the second case is similar. ∎
5. Proof of main theorem
Now we prove the main theorem.
Theorem 5.1.
The full subcategories $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ of the Fukaya category $H^{*}FS(Y^{0},W_{0})$ are equivalent to $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ and $D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, respectively. Furthermore, for the full subcategory $\mathcal{F}:=\mathcal{F}_{1}\cup\mathcal{F}_{2}$, there is an equivalence
(5.1) $\mathcal{F}\xrightarrow[]{\cong}D^{b}_{\mathcal{L}}\mathrm{Coh}(X).$
In particular, the semi-orthogonality in $D^{b}_{\mathcal{L}}Coh(X)$ is respected.
Theorem 5.2 ([Can20, §6]).
There exists a commutative diagram as follows.
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}\mathrm{FS}(Y,-v_{0})$$\scriptstyle{\iota^{*}}$$\scriptstyle{\text{HMS on }V}$$\scriptstyle{\text{HMS on }H=\Sigma_{2}}\scriptstyle{\cup}$
Corollary 5.3 (HMS on $\mathcal{F}_{1}$ factor).
$\mathcal{F}_{1}\cong D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ under $L_{k}\mapsto\mathcal{L}^{k}|_{H}$.
Proof.
By the right vertical arrow in Theorem 5.2, $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)\cong\mathcal{F}_{1}^{\prime}$ for $\mathcal{F}_{1}^{\prime}\subset H^{0}\text{FS}(Y,-v_{0})$ the set of U-shape Lagrangians in $v_{0}:Y\to\mathbb{C}$. Then $\mathcal{F}_{1}^{\prime}\cong\mathcal{F}_{1}$ under the affine transformation on the base from $v_{0}$ to $W_{0}=T^{\epsilon}(v_{0}-1)$. ∎
Corollary 5.4 (HMS on $\mathcal{F}_{2}$ factor).
$\mathcal{F}_{2}\cong D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$ under $K_{j}\mapsto\text{pr}_{V}^{*}\mathcal{L}^{\otimes j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}}$.
Proof.
This follows from Lemma 3.25, HMS for abelian varieties [Fuk02, ACLL], the Künneth formula for Ext, and that $\text{Ext}(\mathcal{O}_{\mathbb{C}},\mathcal{O}_{\mathbb{C}})=\mathbb{C}[x]$. ∎
Proof of Main Theorem 1.8.
The map
$\mathcal{F}=\langle\mathcal{F}_{1},\mathcal{F}_{2}\rangle\to D^{b}_{\mathcal{L}}\mathrm{Coh}(X)=\langle j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1)),\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})\rangle$
defined on $\mathcal{F}_{1}$ by
(5.2) $L_{k}\mapsto j_{*}(p^{*}\mathcal{L}^{\otimes k}|_{H}\otimes\mathcal{O}_{p}(-1))$
and on $\mathcal{F}_{2}$ by
(5.3) $K_{j}\mapsto\pi^{*}(\text{pr}_{V}^{*}\mathcal{L}^{\otimes j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}}).$
is an equivalence within each factor of the semi-orthogonal decomposition by Corollaries 5.3 and 5.4. Since $j:E\to X$ is a closed immersion,
$\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)))=\operatorname{Ext}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{2}}|_{H}).$
For morphisms between factors it is zero on both sides in one direction by Equation (2.1) on the B-side and Lemma 3.24 on the A-side. That is,
$\operatorname{Ext}(\pi^{*}(\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}}),j_{*}(p^{*}i^{*}(\mathcal{L}^{k}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{p}(-1)))=0=HF(K_{j},L_{k}).$
In the other direction, the equivalence on morphisms reduces to HMS for $H$, as follows. By Equation (2.3)
$\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}}))\cong\operatorname{Ext}(\mathcal{L}^{k}|_{H},\mathcal{L}^{j}|_{H})[-1].$
By Corollary 5.3 and Lemma 3.23 this is isomorphic to
$HF(L_{k},L_{j})[-1]\cong HF(L_{k},K_{j})$
because the grading on $W_{0}(K_{i})$ is $\alpha_{i}=0$ which points in the opposite direction from $W_{0}(L_{i})$ constructed by extending as in Figure 4.
To show it’s a functor it remains to prove it respects composition between the two components. This follows from Lemma 4.3 and HMS for abelian varieties [ACLL, Fuk02].